### v1

# Merge then Realign: Simple and Effective Modality-Incremental Continual Learning for Multimodal LLMs

Abstract
Recent advances in Multimodal Large Language Models (MLLMs) have enhanced their versatility as they integrate a growing number of modalities. Considering the heavy cost of training MLLMs, it is necessary to reuse the existing ones and further extend them to more modalities through Modality-incremental Continual Learning (MCL). However, this often comes with a performance degradation in the previously learned modalities. In this work, we revisit the MCL and investigate a more severe issue it faces in contrast to traditional continual learning, that its degradation comes not only from catastrophic forgetting but also from the misalignment between the modality-agnostic and modality-specific components. To address this problem, we propose an elegantly simple MCL paradigm called "MErge then ReAlign" (MERA). Our method avoids introducing heavy training overhead or modifying the model architecture, hence is easy to deploy and highly reusable in the MLLM community. Extensive experiments demonstrate that, despite the simplicity of MERA, it shows impressive performance, holding up to a 99.84% Backward Relative Gain when extending to four modalities, achieving a nearly lossless MCL performance.
Merge then Realign: Simple and Effective Modality-Incremental Continual Learning for Multimodal LLMs
Dingkun Zhang1, Shuhan Qi1thanks: Corresponding author., Xinyu Xiao1, Kehai Chen1, Xuan Wang1, 1Harbin Institute of Technology, Shenzhen dingkunzhang0xffff@gmail.com, {shuhanqi,wangxuan}@cs.hitsz.edu.cn, 23b951021@stu.hit.edu.cn, chenkehai@hit.edu.cn
1 Introduction
With the recent trend of developing general-purpose any-modality Multimodal Large Language Models (MLLMs)(Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024), MLLMs are evolving towards integrating more modalities. The typical MLLM architecture includes modality-specific encoders, modality-specific connectors, and a shared Large Language Model (LLM). A standard process of training MLLMs involves aligning modality-specific components with LLM through modality-text paired data and then fine-tuning on modality-text instruction data. Such architecture and training strategy have been successfully applied to a wide range of modalities, i.e., image(Liu et al., 2024b, a), video(Lin et al., 2024a; Maaz et al., 2024), audio(Li et al., 2024; Wu et al., 2024), point cloud(Chen et al., 2024), etc, equipping MLLMs with the ability to understand a growing number of modalities.
Existing methods(Wu et al., 2024; Zhan et al., 2024; Panagopoulou et al., 2023; Fu et al., 2024) typically employ a joint training strategy, where the MLLM is simultaneously trained on datasets of all predefined modalities. However, it is challenging to extend an existing MLLM to new modalities.
Accordingly, Continual Learning (CL) is proposed to learn from a stream of data. When incrementally learning on the new data, a phenomenon called catastrophic forgetting(McCloskey and Cohen, 1989; Goodfellow et al., 2014) often occurs, i.e., the model forgets the previously learned knowledge. To this end, many CL methods(Kirkpatrick et al., 2017; Yu et al., 2024b; Scialom et al., 2022; Wang et al., 2024) have been proposed to alleviate catastrophic forgetting. Moreover, Modality-incremental Continual Learning (MCL)(Yu et al., 2024a) focuses on the particular scenario of incrementally extending MLLMs to new modalities. However, it is worth noting that: the performance degradation encountered in MCL comes not only from forgetting but also from the misalignment between modality-agnostic and modality-specific components.
To address both forgetting and misalignment, we propose a simple yet effective two-stage MCL paradigm called "MErge then ReAlign" (MERA).
The first stage of MERA aims at addressing the forgetting problem. Inspired by the great success of model merging technique in multi-task learning(Yadav et al., 2024a; Yu et al., 2024c; Yang et al., 2024b, a), we introduce model merging to our MCL framework for the purpose of mitigating forgetting. In this work, we focus on the simplest model merging method, i.e., weight averaging, and revise it into an MCL form, aiming to provide a basic framework. We achieve this by associating its merging coefficients with the progress of CL stages and only merging the modality-agnostic components.
The second stage of MERA aims at addressing the misalignment problem. We leverage a small subset of data from each learned modality to realign the modality encoders with the LLM backbone. In this stage, modality encoders and LLM backbone are both frozen, only the connectors are updated to enable an efficient realignment between them. Further experiments show that the realigning stage can significantly narrow the gap between the incrementally learned MLLM and the individually trained expert MLLMs on each modality.
In summary, the contributions of this paper are threefold:
  • We revisit the Modality-incremental Continual Learning (MCL) and investigate a more severe issue it faces in contrast to traditional continual learning: its performance degradation comes not only from forgetting but also from misalignment.
  • We propose "MErge then ReAlign" (MERA), an elegantly simple and effective two-stage MCL paradigm, to address both forgetting and misalignment.
  • Extensive experiments demonstrate that our MERA significantly outperforms other representative continual learning methods, and even achieves a nearly lossless MCL performance.
Extra-Train-Memory-Free Arch-Modification-Free Replay-Data-Free
Regularization-Based
Architecture-Based $\bullet$
Replay-Based
MERA (Ours)
Table 1: Characteristics of different CL categories and our proposed MERA. Extra-train-memory-free: whether introduces extra GPU memory overhead at training time. Arch-modification-free: whether requires modifying the architecture of the model or adding auxiliary components. Replay-data-free: whether requires access to partial data from the previous tasks or distributions. $\bullet$ denotes that some methods of this category don’t satisfy the property. denotes that this drawback is relatively minor in real applications.
2 Related Work
2.1 Multimodal Large Language Models
Recent advances(Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024) in Multimodal Large Language Models (MLLMs) have extended Large Language Models (LLMs) to perceive multimodal inputs such as image, video, audio, point cloud, etc. Early attempts like Flamingo(Alayrac et al., 2022) integrate vision encoders with LLMs via cross-attention to perform image captioning and visual question answering (VQA) tasks. Subsequent methods like InstructBLIP(Dai et al., 2023) leverage instruction tuning to build a general-purpose multimodal system. These attempts lead to the rapid development of MLLMs.
Among these MLLMs, the most influential one is LLaVA(Liu et al., 2024b, a), which utilizes a simple MLP connector to project visual information encoded by the pre-trained vision encoder into the language embedding space while leveraging visual instruction tuning. Due to its simplicity and effectiveness, LLaVA-like architecture is widely adopted by a wide range of subsequent MLLMs(Lin et al., 2024b, a; Maaz et al., 2024; Wu et al., 2024; Chen et al., 2024). In this paper, we assume that the MLLM has a LLaVA-like architecture that includes modality-specific encoders and connectors, and a shared modality-agnostic LLM.
2.2 Continual Learning
Extending existing MLLMs to new modalities can be viewed as a Continual Learning (CL) problem. CL approaches allow models to continually acquire new knowledge with minor forgetting of previously learned knowledge. Existing CL methods mainly fall into the following three categories: Regularization-based methods(Kirkpatrick et al., 2017; Huszár, 2017; Schwarz et al., 2018) seek to protect the parameters that store important knowledge. However, storing the importance matrix requires extra memory with the same scale of the trainable parameters at training time. Architecture-based methods(Yu et al., 2024a, b; Zadouri et al., 2024) add task-specific parameters to the base model for each new task. The drawbacks of this category are that it requires modifications of the model architecture, harming its reusability, and the model scale often grows linearly as tasks increase, introducing extra memory overhead. Replay-based methods(Scialom et al., 2022; Wang et al., 2024) leverage a small subset of previous data and replay them while learning on new data. A drawback of this category is that it requires access to partial data from the previous tasks or distributions. However, this drawback is relatively minor in real applications as the replay data are often accessible.
Table 1 summarizes the characteristics of different CL categories. Among all the CL methods, replay-based methods are the simplest and most widely used for LLMs since they are extra-train-memory-free and arch-modification-free. In this work, we propose a simple MCL paradigm that has the same advantages as replay-based methods.
2.3 Model Merging
Model merging is an emerging technique often used for multi-task learning(Yadav et al., 2024a; Yu et al., 2024c; Yang et al., 2024b, a). It aims to combine the strengths of multiple isomorphic models into one unified model. The simplest model merging method is weight averaging(Utans, 1996), where weights from each model are averaged to form the merged model. Model merging is based on the assumption that the models being merged lie in the same flat loss basin(Neyshabur et al., 2020), hence their interpolations also have low losses. Aside from its applications in multi-task learning, model merging can be naturally leveraged to reduce forgetting(Marczak et al., 2024).
3 Analysis of Degradations in MCL
Modality-incremental Continual Learning (MCL) is a special scenario of CL, where models incrementally learn on the data from new modalities. The first attempt of MCL is (Yu et al., 2024a). However, we point out that the MCL faces a more severe problem. In contrast to traditional CL scenarios, where the degradation comes solely from the forgetting of old knowledge, the degradation encountered in MCL comes from two aspects: forgetting and misalignment.
Forgetting: the modality-agnostic components (the LLM backbone for MLLM) forget the knowledge of old modalities.
Misalignment: the modality-agnostic components are misaligned with the modality-specific ones. When adapting to a new modality, the LLM backbone and the newly added modality encoder and connector are updated while the old ones are kept frozen. Therefore, the LLM backbone inevitably drifts away from the original multimodal feature space of old modality encoders (even the feature space of the new modality encoder, if its connector is not synchronously updated with LLM backbone). This process results in misalignment, further worsening the degradation. A potential exception to misalignment is when replay-based methods are applied since the old modality-specific components are also updated synchronously. However, empirical results in Section 5.5 suggest that replay-based methods still suffer from a certain degree of misalignment and it can be compensated by our proposed realigning stage.
Due to the existence of misalignment, MCL problem requires special treatments compared to traditional CL problems.
Figure 1: Pipeline of the proposed MERA. The procedures in gray boxes involve training. Refer to caption and Refer to caption represent the frozen and trainable modules, respectively.
4 Method
First, we define the Modality-incremental Continual Learning (MCL) problem as follows. Given a sequence of $m$ modalities $\{M_{1},M_{2},\dots,M_{m}\}$ and their corresponding datasets $\{D_{1},D_{2},\dots,D_{m}\}$, the model sequentially learns on each $D_{i}$ with the permission to access a $r\%$ size subset of previous data as replay data [^1] Different from the definition of replay data in the context of replay-based methods, where $R_{i}$ is sampled from $\{D_{1},D_{2},\dots,D_{i-1}\}$. In this paper, if not explained, $R_{i}$ for replay-based methods is sampled from $\{D_{1},D_{2},\dots,D_{i-1}\}$ while $R_{i}$ for our realigning stage is sampled from $\{D_{1},D_{2},\dots,D_{i}\}$. $R_{i}\leftarrow$ sample $r\%$ data from $\{D_{1},D_{2},\dots,D_{i}\}$. We denote the model after the $i$-th incremental training stage as $\theta_{i}=\{\theta_{i}^{agn},\theta_{i}^{spec}\}$, where $\theta_{i}^{agn},\theta_{i}^{spec}$ denote modality-agnostic and modality-specific components, respectively. For MLLM, the modality-agnostic component is the LLM backbone, and modality-specific components include encoders and connectors of each modality.
To mitigate performance degradations in the previously learned modalities, we propose a two-stage MCL paradigm called "MErge then ReAlign" (MERA). In each stage of MCL, MERA executes the following two stages: merging and realigning, to address the forgetting and misalignment respectively. The overall pipeline of MERA is illustrated in Figure 1.
4.1 Stage 1: Merging
Motivation. Model merging has achieved great success in the field of multi-task learning for its superior capability of combining the strengths of multiple models into one. Moreover, Yadav et al. (2024b) finds that model merging is more effective with larger models. Therefore, applying model merging to large-scale models, such as MLLMs, is inherently beneficial. Inspired by these, we introduce model merging to our MCL framework to mitigate forgetting.
However, model merging methods generally can not be directly applied to continual learning. In this work, we only focus on the simplest model merging method, i.e., weight averaging, and revise it into a CL form to provide a basic framework. To adapt weight averaging to CL, we associate its merging coefficients with the progress of CL stages, forming a Cumulative Moving Average (CMA) merging. At the $i$-th training stage, the CMA merged model is calculated by:
$CMA(\theta_{i-1},\theta_{i};i)=\frac{i-1}{i}\theta_{i-1}+\frac{1}{i}\theta_{i}$
To further adapt to the MCL framework, we only merge the modality-agnostic components (the LLM backbone) and ensemble the modality-specific components (encoders and connectors of each modality):
[…]
$\theta_{i-1}^{spec},\theta_{i}^{spec}\}$
To perform MCL, when starting from $\theta_{i-1}$ to incrementally learn a new modality $M_{i}$, firstly we directly train the model on $D_{i}$ to get $\theta_{i,vanilla}$. Note that the training process here is the same as the standard way of training an MLLM on a single modality, except that the initial weights of the model are inherited from $\theta_{i-1}$. Then secondly, we apply CMA merging to $\theta_{i,vanilla}$ and $\theta_{i-1}$ to derive the merged model $\theta_{i,merged}=CMA^{*}(\theta_{i-1},\theta_{i,vanilla};i)$ with integrated knowledge of both new modality and old modalities.
Discussion. Different from other CL methods, the LLM backbone is asynchronously updated with the new connector in the merging stage. To be specific, in the standard MLLM training step, they are synchronously updated indeed, however, in the CMA merging step, only the LLM backbone is updated, resulting in an additional misalignment between them [^2] This is the reason we sample $R_{i}$ from $\{D_{1},D_{2},\dots,D_{i}\}$ rather than $\{D_{1},D_{2},\dots,D_{i-1}\}$ in our realigning stage..
4.2 Stage 2: Realigning
To address the misalignment issue, we simply leverage a small replay dataset $R_{i}\leftarrow$ sample $r\%$ data from $\{D_{1},D_{2},\dots,D_{i}\}$ to further fine-tune all the connectors of $\theta_{i,merged}$ to obtain the final $\theta_{i}$. This process realigns the encoders of each modality with the LLM backbone. The training objective for realigning is unchanged from the original MLLM training objective, i.e., the auto-regressive loss.
The realigning stage is similar to replay-based CL methods(Scialom et al., 2022; Wang et al., 2024) in form as they both leverage a replay dataset, however, they are essentially different. First, replay methods train on the joint dataset of $D_{i}$ and $R_{i}$, while our realigning stage trains solely on $R_{i}$. Second, the purpose of replay is to review the previous knowledge, while the motivation of realigning stage is to align modality spaces. Third, replay methods update all the parameters to consolidate old knowledge, in contrast, keeping the LLM backbone frozen is crucial in the realigning stage, otherwise the LLM would overfit on replay data.
5 Experiments
5.1 Experimental Setup
We build our MCL experiments on four modalities: image, video, audio, and point cloud, with two different training orders. Based on the prevalence of different modalities, we determine the two orders as follows. Sequential Order: image $\rightarrow$ video $\rightarrow$ audio $\rightarrow$ point cloud. Reverse Order: point cloud $\rightarrow$ audio $\rightarrow$ video $\rightarrow$ image. On top of this, the adopted datasets, metrics, models, and baselines are listed as follows.
Datasets. For each modality $M_{j}$, we leverage a dataset of Captioning (Cap) task and a dataset of Question Answering (QA) task to form the joint dataset $D_{j}=\{D_{j,Cap},D_{j,QA}\}$. The Cap and QA datasets for each modality are listed respectively. For image modality, we use MSCOCO-2014(Lin et al., 2014) and OK-VQA(Marino et al., 2019). For video modality, we use MSVD(Chen and Dolan, 2011) and MSVD-QA(Xu et al., 2017). For audio modality, we use AudioCaps(Kim et al., 2019) and Clotho-AQA(Lipping et al., 2022). For point cloud modality, we use a subset of Cap3D(Luo et al., 2024) and a subset of Cap3D-QA(Panagopoulou et al., 2023). More details of these datasets are in Appendix A.
Evaluation Metrics. First, we leverage Relative Gain(Scialom et al., 2022; Wang et al., 2024) as a normalized metric across different tasks. We train expert MLLMs individually on each single modality $M_{j}$ and test with their respective holdout data, taking their scores on the $k$-th dataset $D_{j,k}$ as upper bound $S_{j,k}^{sup}$. In the incremental stage $i$, the Relative Gain of modality $M_{i}$ with its dataset $D_{j}=\{D_{j,k}\}_{k=1}^{K}$ is calculated by:
$\mathrm{Relative\ Gain}_{j}^{i}=\frac{1}{K}\sum_{k=1}^{K}\frac{S_{j,k}^{i}}{S_{j,k}^{sup}}$
where $S_{j,k}^{i}$ is the score on the test set of $D_{j,k}$ in the stage $i$. Here, we utilize CIDEr score(Vedantam et al., 2015) and prediction accuracy (Acc) for Cap and QA tasks respectively to calculate $S_{j,k}^{sup}$ and $S_{j,k}^{i}$. To evaluate the performance degradations of the previously learned modalities, we calculate the Backward Relative Gain in the stage $i$ as:
$\mathrm{Bw\ Relative\ Gain}^{i}=\frac{1}{i-1}\sum_{j=1}^{i-1}\mathrm{Relative\ Gain}_{j}^{i}$
To measure the plasticity, i.e., the ability to adapt to new knowledge, we calculate the Forward Relative Gain in the stage $i$ as:
$\mathrm{Fw\ Relative\ Gain}^{i}=\mathrm{Relative\ Gain}_{i}^{i}$
Model and Training Details. We leverage the mainstream MLLM architecture, i.e., LLaVA-like architecture with the Llama-3-8B-Instruct (Dubey et al., 2024) as its LLM backbone. The selections of modality encoders and connectors are detailed in Appendix B. Trainings that involve updating the LLM backbone utilize LoRA(Hu et al., 2022) for parameter-efficient fine-tuning. The training process in our merging stage is the same as the regular MLLM training, i.e., in the first step, only the connector is updated with Cap datasets, then in the second step, the connector and the LLM backbone are updated with all the task-related datasets (the combination of Cap and QA datasets in our case). In our realigning stage, the replay datasets are randomly sampled from the joint datasets of Cap and QA tasks. For each training process, the hyperparameters are listed in Appendix B.
Figure 2: Progressive Backward Relative Gain in modality-incremental continual learning. For each stage $i$, we plot the average score of corresponding Backward Relative Gain with two different training orders. We set Backward Relative Gain to 100% for the $1$-st stage, denoting the initial performance without degradation.
Baselines. In our experiments, we compare our MERA with non-CL fine-tuning and the representative methods of each CL category: Fine-Tuning: directly train MLLMs sequentially on each modality without applying any CL method. Replay: the vanilla replay-based CL method. During training on a new task, the model is updated with both samples from the current task and a set of randomly sampled replay data from previous tasks, to review knowledge of earlier tasks while learning the new one. EWC(Kirkpatrick et al., 2017): the most representative regularization-based CL method. EWC mitigates forgetting by restricting the updates of important weights during training on new tasks. It uses the Fisher information matrix to measure the importance of each weight. PathWeave(Yu et al., 2024a): an architecture-based CL method, also the first MCL method for MLLMs. PathWeave uses an adapter-in-adapter mechanism to memorize and extract knowledge from historical modalities to enhance the learning of the current modality. PathWeave is originally built on X-InstructBLIP(Panagopoulou et al., 2023). For a fair comparison, we implement PathWeave for our adopted MLLM architecture [^3] In their original settings, PathWeave removes the newly added modules when testing the former modalities. However, this results in the inability to perform cross-modality tasks, which are common in real applications. Therefore, we do not remove them for a fair comparison.. Implementation details of each baseline method are in Appendix D.
Method Sequential Reverse
Mean Std Mean Std
Fine-Tuning 59.76 27.23 48.96 35.70
Replay (1%) 66.09 25.86 43.95 39.03
Replay (10%) 77.52 16.32 51.90 36.34
EWC 74.93 17.14 68.01 21.54
PathWeave 86.85 12.17 80.09 13.31
MERA (1%) 97.90 6.02 84.42 12.93
MERA (10%) 101.00 3.90 93.42 6.25
Table 2: The mean and standard deviation of Backward Relative Gains in all the training stages. Results are reported on different training orders. The best results are in bold, while the second-best are underlined.
Method Training Inference
Peak Memory Time-Consuming Peak Memory Lantency per Token
Fine-Tuning 37.43 GB 53 h 17.71 GB 34 ms
Replay (1%) 37.43 GB 54 h 17.71 GB 34 ms
Replay (10%) 37.43 GB 59 h 17.71 GB 34 ms
EWC 38.73 GB 54 h 17.71 GB 34 ms
PathWeave 40.08 GB 81 h 20.32 GB 111 ms
MERA (1%) 37.43 GB 54 h 17.71 GB 34 ms
MERA (10%) 37.43 GB 61 h 17.71 GB 34 ms
Table 3: Training and inference overheads of different methods. The peak memories during training and inference are measured with batch sizes of 4 and 1 respectively. The time-consuming refers to the total GPU hours for continually learning the four modalities. All metrics are measured on a single NVIDIA RTX A6000 48G. The non-optimal results are colored in red.
Method Sequential Reverse
Mean Std Mean Std
Fine-Tuning 59.76 27.23 48.96 35.70
+Merging 90.29 7.42 70.00 30.87
+Realigning 87.92 12.41 71.90 24.52
MERA 101.00 3.90 93.42 6.25
Table 4: Ablation study of different components in MERA. The realigning stage uses 10% replay data. The mean and standard deviation of Backward Relative Gains in all the training stages are reported on different training orders. The best results are in bold.
5.2 Main Results
We conduct experiments under our MCL setting with both sequential and reverse orders. For Replay and our MERA, results using $r\%,r=\{1,10\}$ replay data are reported, denoted by Replay ($r\%$) and MERA ($r\%$) respectively. The progressive Backward Relative Gains averaged from different training orders are plotted in Figure 2. It is observed that our MERA demonstrates an impressive capability of mitigating performance degradation with consistent and promising Backward Relative Gains. When extending to all the four modalities, MERA ($10\%$) holds up to a 99.84% Backward Relative Gain, indicating that MERA can achieve a nearly lossless MCL performance, with at least 20.37% absolute improvements of Backward Relative Gain compared with other baselines. Notably, when only leveraging $1\%$ replay data, MERA (1%) can still achieve at least 15.94% absolute improvements over other baselines.
Further, we calculated the mean and standard deviation of Backward Relative Gains in all training stages for each method, in different training orders. Table 2 shows that our MERA (10%) achieves the highest mean and lowest standard deviation in both training orders, indicating its superior performance and high stability. Notably, in sequential order, MERA (10%) performs even better than lossless MCL, with an over 100% average Backward Relative Gain, also at least 14.15% absolute improvements of average Backward Relative Gain over other baselines. In reverse order, MERA (10%) also achieves at least 13.33% absolute improvements. When with only 1% replay data, MERA (1%) still achieves at least 11.05% and 4.33% absolute improvements in sequential and reverse orders respectively.
From Figure 2 and its raw data shown in Appendix E, we also observe a faint phenomenon of Positive Backward Transfer(Lin et al., 2022) that learning new knowledge improves the performance on previously learned tasks. For most CL methods and non-CL fine-tuning, the Backward Relative Gain comes to a low level when incrementally learning the first modality (corresponding to the $2$-nd training stage), but starts to stabilize or even increase when incrementally learning more modalities.
5.3 Efficiency Comparisons
We further compare the efficiency of different baselines and our MERA, as shown in Table 3. It is observed that our MERA can achieve optimal results except that MERA (1%) and MERA (10%) introduce 2% and 15% extra training time-consuming respectively. However, we believe its trade-off between training time-consuming and performance is worthwhile, considering the impressive performance of MERA. In our experiments, EWC and PathWeave introduce marginal extra training memory overhead, as we employ parameter-efficient fine-tuning. However, for larger LoRA ranks or even full model fine-tuning, their extra training memory consumptions would be substantial, as they necessitate storing additional parameters whose sizes increase linearly with the trainable parameters.
5.4 Ablation Study
We conduct ablation studies to investigate the effectiveness of each component in MERA. Results are shown in Table 4. Firstly, from Table 4 and Table 2, it is observed that the merging stage alone can already beat many other baselines, achieving the best and second-best performances among baselines in sequential and reverse orders respectively. Secondly, the realigning stage alone can also achieve performance improvements by addressing the misalignment issue. Combining both the merging and realigning stages, MERA further narrows the gap between the incrementally learned models and the individually trained experts on each modality, even, surpassing the individually trained experts in sequential training order with over 100% Backward Relative Gain.
Method Sequential Reverse
Mean Std Mean Std
Fine-Tuning 59.76 27.23 48.96 35.70
+Realigning +28.16 -14.83 +22.93 -11.17
Replay (1%) 66.09 25.86 43.95 39.03
+Realigning +20.64 -9.75 +24.35 -8.36
Replay (10%) 77.52 16.32 51.90 36.34
+Realigning +14.21 -6.67 +23.71 -11.07
EWC 74.93 17.14 68.01 21.54
+Realigning +19.54 -6.87 +23.02 -13.25
PathWeave 86.85 12.17 80.09 13.31
+Realigning +6.22 -0.28 +2.40 -1.53
Merging 90.29 7.42 70.00 30.87
+Realigning +10.71 -3.52 +23.42 -24.63
Table 5: Applying realigning to different CL or non-CL methods can further improve their Backward Relative Gain and stability. The realigning stage uses 10% replay data. The increased mean indicates better performance, while the decreased standard deviation indicates better stability. Improvements are colored in green.
Figure 3: Progressive Forward Relative Gain in modality-incremental continual learning. For each stage $i$, we plot the average score of corresponding Forward Relative Gain with two different training orders. We set Forward Relative Gain to 100% for the $1$-st stage, denoting the initial lossless plasticity.
5.5 Misalignment Is Common in MCL
Since the realigning stage achieves great success on top of our proposed merging stage, we further ask another question: does realigning benefit other CL methods, or is misalignment a common phenomenon in MCL? To examine this, we perform the realigning stage at the end of every training stage for different CL or non-CL methods to observe whether there are performance improvements. Table 5 shows that the additional realigning stage brings substantial performance improvements and increased stability for different CL or non-CL methods. Based on this observation, we can conclude that misalignment is a common phenomenon in MCL, and can be compensated by our proposed realigning stage.
5.6 Plasticity Analysis
Aside from alleviating performance degradation, the capability to adapt to new knowledge, i.e., plasticity, is also an important aspect. We use Forward Relative Gain as the metric to evaluate it. The progressive Forward Relative Gains averaged from different training orders are plotted in Figure 3. It is observed that the most elastic CL methods are EWC and PathWeave, while our MERA (10%) demonstrates comparable plasticity.
Notably, from Figure 3, we observe a strong phenomenon of Positive Forward Transfer(Ke et al., 2021) that the knowledge or skills acquired from earlier tasks improve the learning efficiency of new tasks. The Positive Forward Transfer emerges before the $4$-th training stage, on EWC, PathWeave, and MERA. This phenomenon is also reported by other MCL literature(Yu et al., 2024a). In contrast to Positive Forward Transfer, there is a gradual plasticity loss(Dohare et al., 2024, 2023) as the model attempts to retain more knowledge. This explains the decreases in Forward Relative Gain across different CL methods in the $4$-th stage, as the plasticity loss comes to a dominant position.
6 Future Work
This work is one of the early attempts of MCL, mainly focusing on designing effective methods and addressing the misalignment issue. However, our extensive experiments suggest that there are many other interesting aspects of MCL that warrant attention. In Section 5.2 and Section 5.6, we observe the phenomenons of Positive Backward Transfer and Positive Forward Transfer, respectively. This could be related to the complex cross-modal interaction in multimodal learning, urging for further research on the mechanisms of modality interaction and methods that boost a positive modality interaction, in the context of MCL.
7 Conclusion
In this paper, we mainly focus on designing effective methods for Modality-incremental Continual Learning (MCL). First, we revisit MCL and investigate a more severe issue it faces in contrast to traditional continual learning that its performance degradation comes not only from forgetting but also from misalignment. To address both the forgetting and misalignment, we propose MERA, a simple yet effective MCL paradigm. Extensive experiments demonstrate that MERA significantly outperforms the baselines, and even achieves a nearly lossless MCL performance. Further, we observe a sign of complex cross-modal interaction in MCL, providing a direction for future work.
Limitations
Our work is restricted in the following aspects. First, our experiments are limited to four commonly used modalities due to the lack of resources for other less-studied modalities. Second, our experiments are limited to two types of tasks for each modality due to the lack of general-purpose and diverse instruction datasets and their corresponding benchmarks for some modalities, i.e., audio and point cloud. Third, our method is limited to any-to-text MLLMs while there is now a trend of exploring any-to-any MLLMs. However, the main idea of MERA is generic to any-to-any MLLMs since their architecture topologies are similar. Fourth, our work only explores the simplest model merging method in the context of MCL, aiming to provide a universal framework, leaving the adaptation of other model merging methods to MCL for future work.
Appendix
Appendix A Dataset Details
Table 6 details the statistics of each dataset. Some datasets are filtered from their original ones:
  • MSCOCO-2014(Lin et al., 2014): Each image has multiple captions, we only use its first caption to form the training set.
  • Clotho-AQA(Lipping et al., 2022): Each sample is annotated with a confidence level, we only use the samples whose confidence levels are "yes" to form the training set and test set.
  • Cap3D(Luo et al., 2024): Since the original dataset is huge in scale, we filter out the samples whose caption is longer than 100 letters. Then, we randomly sample a 50K subset as the training set.
  • Cap3D-QA(Panagopoulou et al., 2023): Since the original dataset is huge in scale, we randomly sample a 30K subset as the training set.
For each dataset, we use a randomly sampled 1K subset of its holdout test set as the final test set, except for the MSVD(Chen and Dolan, 2011), since the size of its original test set is less than 1K.
Appendix B Implementation Details
We build our experimental codebase on top of LLaVA(Liu et al., 2024b, a) and NExT-GPT(Wu et al., 2024). We detail the modality-specific components of each modality as follows:
  • Image: We use CLIP-ViT-L-336px(Openai, 2021) as the pre-trained image encoder, a randomly initialized MLP as the connector.
  • Video: We use CLIP-ViT-L-336px(Openai, 2021) as the pre-trained video encoder, a randomly initialized MLP as the connector. We uniformly sample 4 frames from a video as input frames. Then each frame is encoded by the video encoder separately. The output feature frames are downsampled by 2x using bilinear pooling before sending into the connector to improve efficiency.
[…]
  • Point Cloud: We use Point-BERT-v1.2Xu et al. (2024) as the pre-trained point cloud encoder, a randomly initialized MLP as the connector.
We set the hyperparameters mainly following previous worksLiu et al., 2024b, a, as listed in Table 7. For training that involves updating the LLM backbone, we utilize parameter-efficient fine-tuning with LoRA(Hu et al., 2022) applied across all linear modules within the LLM, setting the LoRA rank to 128 and the alpha parameter to 128. All the experiments are conducted on a single NVIDIA RTX A6000 48G with FP16.
#Training Set #Test Set License
MSCOCO-2014* 82K 1K CC-BY 4.0
OK-VQA 26K 1K CC-BY 4.0
MSVD 48K 670 -
MSVD-QA 30K 1K -
AudioCaps 44K 1K -
Clotho-AQA* 15K 1K MIT License
Cap3D* 50K 1K ODC-BY 1.0
Cap3D-QA* 30K 1K -
Table 6: Statistics of the datasets. Datasets marked with * are filtered from their original ones.
Hyperparameters Std-Stage 1 Std-Stage 2 Realigning
Trainable Components Connectors LLM and Connectors Connectors
Batch Size 128 16 16
Learning Rate of Connectors 1e-3 2e-5 2e-5
Learning Rate of LLM - 2e-4 -
Learning Rate Schedule Cosine Decay
Warmup Ratio 0.03
Epoch 1
Table 7: Hyperparameters for each training stage. Std-stage 1 and Std-stage 2 refer to the stage 1 and 2 of the standard MLLM training process.
Appendix C Evaluation Details
For the calculation of CIDEr scores(Vedantam et al., 2015), we utilized an open-sourced library CaptionMetrics(wangleihitcs, 2019). For the calculation of prediction accuracy, we leverage a GPT-based open-ended QA evaluation with GPT-4o mini as the judge model. The GPT is prompted to judge whether the generated prediction semantically matches the ground truth answer. The prompt template is shown in Table 8. All the reported experimental results are from single runs.
System Prompt:
You are an intelligent chatbot designed for evaluating the correctness of generative outputs for question-answer pairs. Your task is to compare the predicted answer with the correct answer and determine if they match meaningfully. Here’s how you can accomplish the task:
##INSTRUCTIONS:
- Focus on the meaningful match between the predicted answer and the correct answer.
- Consider synonyms or paraphrases as valid matches.
- Evaluate the correctness of the prediction compared to the answer.
User Prompt:
Please evaluate the following question-answer pair:
Question: <question>
Correct Answer: <answer>
Predicted Answer: <prediction>
Provide your evaluation only as a yes/no and score where the score is an integer value between 0 and 5, with 5 indicating the highest meaningful match. Please generate the response in the form of a Python dictionary string with keys ’pred’ and ’score’, where value of ’pred’ is a string of ’yes’ or ’no’ and value of ’score’ is in INTEGER, not STRING.DO NOT PROVIDE ANY OTHER OUTPUT TEXT OR EXPLANATION. Only provide the Python dictionary string. For example, your response should look like this: {’pred’: ’yes’, ’score’: 4.8}.
Table 8: Prompts to query the GPT for open-ended QA evaluation. The placeholders in red boxes are filled according to each evaluated sample.
Appendix D Implementation of Baselines
The implementation details of each CL baseline are listed as follows:
  • Replay leverage a replay dataset $R_{i}\leftarrow$ sample $r\%$ data from $\{D_{1},D_{2},\dots,D_{i-1}\}$ when training on a new modality $M_{i}$. It can be seen as training on a joint dataset of $R_{i}$ and $D_{i}$.
  • EWC firstly estimates the Fisher information matrix $F_{i-1}$ of the last training stage $i-1$ as:
    $F_{i-1}=$
    $\mathbb{E}_{x\sim D_{i-1}}\nabla_{\theta_{i-1}}\mathcal{L}(\theta_{i-1},x)\cdot\nabla_{\theta_{i-1}}\mathcal{L}(\theta_{i-1},x)^{T}$
    where $\mathcal{L}(\theta_{i-1},x)$ denotes the auto-regressive loss of model $\theta_{i-1}$ on data $x\sim D_{i-1}$, which is sampled from a 1% size random subset of $D_{i-1}$. Then the loss function $\mathcal{L}^{*}(\theta_{i},x)$ of stage $i$ is:
    $\mathcal{L}^{*}(\theta_{i},x)=\mathcal{L}(\theta_{i},x)+\sum_{j=0}^{i-1}\frac{\lambda}{2}F_{j}(\theta_{i}-\theta_{i-1})^{2}$
    where the hyperparameter $\lambda$ is set to 1 as default. This implementation of $\mathcal{L}^{*}(\theta_{i},x)$ is known as Online EWC(Huszár, 2017; Schwarz et al., 2018).
  • PathWeave leverages Adapter-in-Adapter (AnA) modules. In our implementation, the AnA modules are injected in the LLM rather than the connector for a fair comparison. The rank of AnA is consistent with the LoRA rank of other baselines, which is 128.
Appendix E Complete Raw Data
Table 9 and Table 10 show the raw data of Figure 2 and Figure 3.
Method Image Video Audio Point Cloud
MSCOCO OK-VQA MSVD MSVD-QA AudioCaps Clotho-AQA Cap3D Cap3D-QA
Individually Trained Experts 100.76 0.358 138.39 0.460 60.14 0.658 99.93 0.568
Fine-Tuning Stage 1 100.76 0.358 - - - - - -
Stage 2 54.52 0.172 130.22 0.555 - - - -
Stage 3 34.87 0.303 12.78 0.292 43.17 0.590 - -
Stage 4 58.63 0.201 29.55 0.350 8.28 0.094 84.40 0.524
Replay (1%) Stage 1 100.76 0.358 - - - - - -
Stage 2 41.45 0.125 137.07 0.569 - - - -
Stage 3 65.79 0.276 30.74 0.312 55.21 0.675 - -
Stage 4 59.94 0.225 102.16 0.469 22.17 0.490 81.43 0.508
Replay (10%) Stage 1 100.76 0.358 - - - - - -
Stage 2 50.65 0.266 137.67 0.584 - - - -
Stage 3 83.32 0.318 33.87 0.381 44.82 0.651 - -
Stage 4 67.42 0.259 133.43 0.520 24.13 0.525 73.19 0.515
EWC Stage 1 100.76 0.358 - - - - - -
Stage 2 64.84 0.208 155.09 0.595 - - - -
Stage 3 44.22 0.211 73.14 0.569 59.86 0.690 - -
Stage 4 56.54 0.227 36.72 0.564 26.64 0.651 96.40 0.551
PathWeave Stage 1 100.76 0.358 - - - - - -
Stage 2 78.06 0.234 158.51 0.606 - - - -
Stage 3 79.07 0.251 138.63 0.547 59.47 0.682 - -
Stage 4 66.92 0.255 123.39 0.536 38.53 0.639 97.32 0.554
MERA (1%) Stage 1 100.76 0.358 - - - - - -
Stage 2 93.70 0.304 153.73 0.573 - - - -
Stage 3 90.42 0.316 147.42 0.567 57.09 0.678 - -
Stage 4 95.18 0.334 142.67 0.562 53.04 0.678 79.32 0.454
MERA (10%) Stage 1 100.76 0.358 - - - - - -
Stage 2 98.30 0.340 152.20 0.579 - - - -
Stage 3 96.46 0.346 147.89 0.566 61.49 0.684 - -
Stage 4 98.05 0.338 141.25 0.560 56.79 0.678 87.59 0.468
Table 9: Raw data of sequential order training. Results that are better than the last stage are colored in green, indicating a Positive Backward Transfer.
Method Point Cloud Audio Video Image
Cap3D Cap3D-QA AudioCaps Clotho-AQA MSVD MSVD-QA MSCOCO OK-VQA
Individually Trained Experts 99.93 0.568 60.14 0.658 138.39 0.460 100.76 0.358
Fine-Tuning Stage 1 99.93 0.568 - - - - - -
Stage 2 2.74 0.178 39.25 0.519 - - - -
Stage 3 37.26 0.280 21.34 0.158 121.29 0.550 - -
Stage 4 26.69 0.199 23.11 0.519 23.40 0.266 86.12 0.342
Replay (1%) Stage 1 99.93 0.568 - - - - - -
Stage 2 0.98 0.101 48.48 0.640 - - - -
Stage 3 35.18 0.337 8.30 0.138 124.89 0.546 - -
Stage 4 9.39 0.192 17.67 0.498 1.06 0.255 83.38 0.347
Replay (10%) Stage 1 99.93 0.568 - - - - - -
Stage 2 0.63 0.171 47.38 0.641 - - - -
Stage 3 68.26 0.470 12.83 0.372 134.07 0.575 - -
Stage 4 3.42 0.241 18.33 0.540 2.81 0.228 87.31 0.342
EWC Stage 1 99.93 0.568 - - - - - -
Stage 2 29.97 0.442 59.31 0.672 - - - -
Stage 3 19.91 0.375 29.25 0.611 148.26 0.578 - -
Stage 4 45.25 0.327 23.39 0.511 47.53 0.524 98.91 0.320
PathWeave Stage 1 99.93 0.568 - - - - - -
Stage 2 71.79 0.420 55.07 0.648 - - - -
Stage 3 63.75 0.380 38.73 0.628 148.61 0.577 - -
Stage 4 55.23 0.370 37.23 0.603 85.67 0.521 87.04 0.361
MERA (1%) Stage 1 99.93 0.568 - - - - - -
Stage 2 64.79 0.470 58.77 0.684 - - - -
Stage 3 66.76 0.377 43.00 0.595 147.99 0.547 - -
Stage 4 70.40 0.387 51.02 0.651 140.20 0.538 93.33 0.362
MERA (10%) Stage 1 99.93 0.568 - - - - - -
Stage 2 87.10 0.505 58.99 0.695 - - - -
Stage 3 79.24 0.437 58.65 0.650 145.56 0.552 - -
Stage 4 81.05 0.425 60.28 0.653 146.72 0.569 97.62 0.367
Table 10: Raw data of reverse order training. Results that are better than the last stage are colored in green, indicating a Positive Backward Transfer.

### v2

@@
 Recent advances in Multimodal Large Language Models (MLLMs) have enhanced their versatility as they integrate a growing number of modalities.
-Considering the heavy cost of training MLLMs, it is necessary to reuse the existing ones and further extend them to more modalities through Modality-incremental Continual Learning (MCL).
-However, this often comes with a performance degradation in the previously learned modalities.
-In this work, we revisit the MCL and investigate a more severe issue it faces in contrast to traditional continual learning, that its degradation comes not only from catastrophic forgetting but also from the misalignment between the modality-agnostic and modality-specific components.
-To address this problem, we propose an elegantly simple MCL paradigm called "MErge then ReAlign" (MERA).
-Our method avoids introducing heavy training overhead or modifying the model architecture, hence is easy to deploy and highly reusable in the MLLM community.
-Extensive experiments demonstrate that, despite the simplicity of MERA, it shows impressive performance, holding up to a 99.84% Backward Relative Gain when extending to four modalities, achieving a nearly lossless MCL performance.
+Considering the heavy cost of training MLLMs, it is efficient to reuse the existing ones and extend them to more modalities through Modality-incremental Continual Learning (MCL).
+The exploration of MCL is in its early stages.
+In this work, we dive into the causes of performance degradation in MCL.
+We uncover that it suffers not only from forgetting as in traditional continual learning, but also from misalignment between the modality-agnostic and modality-specific components.
+To this end, we propose an elegantly simple MCL paradigm called "MErge then ReAlign" (MERA) to address both forgetting and misalignment.
+MERA avoids introducing heavy model budgets or modifying model architectures, hence is easy to deploy and highly reusable in the MLLM community.
+Extensive experiments demonstrate the impressive performance of MERA, holding an average of 99.84% Backward Relative Gain when extending to four modalities, achieving nearly lossless MCL performance.
+Our findings underscore the misalignment issue in MCL.
+More broadly, our work showcases how to adjust different components of MLLMs during continual learning.
 Merge then Realign: Simple and Effective Modality-Incremental Continual Learning for Multimodal LLMs
-Dingkun Zhang1, Shuhan Qi1thanks: Corresponding author., Xinyu Xiao1, Kehai Chen1, Xuan Wang1, 1Harbin Institute of Technology, Shenzhen dingkunzhang0xffff@gmail.com, {shuhanqi,wangxuan}@cs.hitsz.edu.cn, 23b951021@stu.hit.edu.cn, chenkehai@hit.edu.cn
+Dingkun Zhang1, Shuhan Qi1,2thanks: Corresponding author., Xinyu Xiao1, Kehai Chen1, Xuan Wang1 1Harbin Institute of Technology, Shenzhen 2Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies dingkunzhang0xffff@gmail.com, {shuhanqi,wangxuan}@cs.hitsz.edu.cn, 23b951021@stu.hit.edu.cn, chenkehai@hit.edu.cn
 1 Introduction
-With the recent trend of developing general-purpose any-modality Multimodal Large Language Models (MLLMs)(Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024), MLLMs are evolving towards integrating more modalities.
+With the recent trend of developing general-purpose any-modality Multimodal Large Language Models (MLLMs) (Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024; Shuhan et al., 2023), MLLMs are evolving towards integrating more modalities.
 The typical MLLM architecture includes modality-specific encoders, modality-specific connectors, and a shared Large Language Model (LLM).
-A standard process of training MLLMs involves aligning modality-specific components with LLM through modality-text paired data and then fine-tuning on modality-text instruction data.
-Such architecture and training strategy have been successfully applied to a wide range of modalities, i.e., image(Liu et al., 2024b, a), video(Lin et al., 2024a; Maaz et al., 2024), audio(Li et al., 2024; Wu et al., 2024), point cloud(Chen et al., 2024), etc, equipping MLLMs with the ability to understand a growing number of modalities.
-Existing methods(Wu et al., 2024; Zhan et al., 2024; Panagopoulou et al., 2023; Fu et al., 2024) typically employ a joint training strategy, where the MLLM is simultaneously trained on datasets of all predefined modalities.
-However, it is challenging to extend an existing MLLM to new modalities.
-Accordingly, Continual Learning (CL) is proposed to learn from a stream of data.
-When incrementally learning on the new data, a phenomenon called catastrophic forgetting(McCloskey and Cohen, 1989; Goodfellow et al., 2014) often occurs, i.e., the model forgets the previously learned knowledge.
-To this end, many CL methods(Kirkpatrick et al., 2017; Yu et al., 2024b; Scialom et al., 2022; Wang et al., 2024) have been proposed to alleviate catastrophic forgetting.
-Moreover, Modality-incremental Continual Learning (MCL)(Yu et al., 2024a) focuses on the particular scenario of incrementally extending MLLMs to new modalities.
-However, it is worth noting that: the performance degradation encountered in MCL comes not only from forgetting but also from the misalignment between modality-agnostic and modality-specific components.
+A standard process of training MLLMs involves aligning modality-specific components with LLM through modality-text paired data and then fine-tuning on modality-text instruction data (Rao et al., 2024).
+Such architecture and training strategy have been successfully applied to a wide range of modalities, i.e., image (Liu et al., 2024b, a; Rao et al., 2023), video (Lin et al., 2024a; Maaz et al., 2024), audio (Li et al., 2024; Wu et al., 2024), point cloud (Chen et al., 2024), etc, equipping MLLMs with the ability to understand a growing number of modalities.
+Existing methods (Wu et al., 2024; Zhan et al., 2024; Panagopoulou et al., 2023; Fu et al., 2024) typically employ a joint training strategy, where the MLLM is jointly trained on datasets of all predefined modalities (Xin et al., 2024).
+However, it is challenging to extend an existing MLLM to new modalities as it requires another round of joint training on the previous modalities and the new modalities.
+To reuse the existing models and adapt them to new data, Continual Learning (CL) is proposed to learn from a stream of data.
+During continual learning, performance degradation in previously learned tasks often occurs.
+The degradation is generally attributed to catastrophic forgetting (McCloskey and Cohen, 1989; Goodfellow et al., 2014; Rao et al., 2025), i.e., the model forgets the previously learned knowledge.
+To this end, many CL methods (Kirkpatrick et al., 2017; Yu et al., 2024b; Scialom et al., 2022; Wang et al., 2024b) have been proposed to alleviate catastrophic forgetting.
+In addition to traditional CL, Modality-incremental Continual Learning (MCL) (Yu et al., 2024a) focuses on the particular scenario of incrementally extending MLLMs to new modalities.
+The exploration of MCL is in its early stages.
+In this work, we first analyze the causes of performance degradation in MCL.
+Unlike traditional CL, the performance degradation encountered in MCL comes not only from forgetting but also from the misalignment between modality-agnostic and modality-specific components.
 To address both forgetting and misalignment, we propose a simple yet effective two-stage MCL paradigm called "MErge then ReAlign" (MERA).
 The first stage of MERA aims at addressing the forgetting problem.
-Inspired by the great success of model merging technique in multi-task learning(Yadav et al., 2024a; Yu et al., 2024c; Yang et al., 2024b, a), we introduce model merging to our MCL framework for the purpose of mitigating forgetting.
-In this work, we focus on the simplest model merging method, i.e., weight averaging, and revise it into an MCL form, aiming to provide a basic framework.
+To this end, we introduce model merging to our MCL framework.
+We focus on the simplest model merging method, i.e., weight averaging, and revise it into an MCL form.
 We achieve this by associating its merging coefficients with the progress of CL stages and only merging the modality-agnostic components.
@@
 We leverage a small subset of data from each learned modality to realign the modality encoders with the LLM backbone.
-In this stage, modality encoders and LLM backbone are both frozen, only the connectors are updated to enable an efficient realignment between them.
+In this stage, modality encoders and LLM backbone are both frozen, only the lightweight connectors are updated to enable an efficient realignment between them.
 Further experiments show that the realigning stage can significantly narrow the gap between the incrementally learned MLLM and the individually trained expert MLLMs on each modality.
 In summary, the contributions of this paper are threefold:
-  • We revisit the Modality-incremental Continual Learning (MCL) and investigate a more severe issue it faces in contrast to traditional continual learning: its performance degradation comes not only from forgetting but also from misalignment.
+  • We analyze the causes of degradation in Modality-incremental Continual Learning (MCL).
+We uncover that it suffers not only from forgetting as in traditional continual learning, but also from misalignment.
   • We propose "MErge then ReAlign" (MERA), an elegantly simple and effective two-stage MCL paradigm, to address both forgetting and misalignment.
-  • Extensive experiments demonstrate that our MERA significantly outperforms other representative continual learning methods, and even achieves a nearly lossless MCL performance.
+  • Extensive experiments show that our MERA significantly outperforms other representative continual learning methods including the state-of-the-art MCL method, and even achieves nearly lossless MCL performance.
 Extra-Train-Memory-Free Arch-Modification-Free Replay-Data-Free
@@
 Replay-Based
+Merging-Based
 MERA (Ours)
 Table 1: Characteristics of different CL categories and our proposed MERA.
-Extra-train-memory-free: whether introduces extra GPU memory overhead at training time.
-Arch-modification-free: whether requires modifying the architecture of the model or adding auxiliary components.
-Replay-data-free: whether requires access to partial data from the previous tasks or distributions.
+Extra-train-memory-free: does not introduce extra GPU memory overhead at training time.
+Arch-modification-free: does not modify the architecture of the model or add auxiliary components.
+Replay-data-free: does not require access to partial data from the previous tasks or distributions.
 $\bullet$ denotes that some methods of this category don’t satisfy the property.
@@
 2.1 Multimodal Large Language Models
-Recent advances(Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024) in Multimodal Large Language Models (MLLMs) have extended Large Language Models (LLMs) to perceive multimodal inputs such as image, video, audio, point cloud, etc.
-Early attempts like Flamingo(Alayrac et al., 2022) integrate vision encoders with LLMs via cross-attention to perform image captioning and visual question answering (VQA) tasks.
-Subsequent methods like InstructBLIP(Dai et al., 2023) leverage instruction tuning to build a general-purpose multimodal system.
-These attempts lead to the rapid development of MLLMs.
-Among these MLLMs, the most influential one is LLaVA(Liu et al., 2024b, a), which utilizes a simple MLP connector to project visual information encoded by the pre-trained vision encoder into the language embedding space while leveraging visual instruction tuning.
-Due to its simplicity and effectiveness, LLaVA-like architecture is widely adopted by a wide range of subsequent MLLMs(Lin et al., 2024b, a; Maaz et al., 2024; Wu et al., 2024; Chen et al., 2024).
-In this paper, we assume that the MLLM has a LLaVA-like architecture that includes modality-specific encoders and connectors, and a shared modality-agnostic LLM.
-2.2 Continual Learning
-Extending existing MLLMs to new modalities can be viewed as a Continual Learning (CL) problem.
-CL approaches allow models to continually acquire new knowledge with minor forgetting of previously learned knowledge.
-Existing CL methods mainly fall into the following three categories: Regularization-based methods(Kirkpatrick et al., 2017; Huszár, 2017; Schwarz et al., 2018) seek to protect the parameters that store important knowledge.
-However, storing the importance matrix requires extra memory with the same scale of the trainable parameters at training time.
-Architecture-based methods(Yu et al., 2024a, b; Zadouri et al., 2024) add task-specific parameters to the base model for each new task.
-The drawbacks of this category are that it requires modifications of the model architecture, harming its reusability, and the model scale often grows linearly as tasks increase, introducing extra memory overhead.
-Replay-based methods(Scialom et al., 2022; Wang et al., 2024) leverage a small subset of previous data and replay them while learning on new data.
-A drawback of this category is that it requires access to partial data from the previous tasks or distributions.
-However, this drawback is relatively minor in real applications as the replay data are often accessible.
-Table 1 summarizes the characteristics of different CL categories.
-Among all the CL methods, replay-based methods are the simplest and most widely used for LLMs since they are extra-train-memory-free and arch-modification-free.
-In this work, we propose a simple MCL paradigm that has the same advantages as replay-based methods.
-2.3 Model Merging
-Model merging is an emerging technique often used for multi-task learning(Yadav et al., 2024a; Yu et al., 2024c; Yang et al., 2024b, a).
-It aims to combine the strengths of multiple isomorphic models into one unified model.
-The simplest model merging method is weight averaging(Utans, 1996), where weights from each model are averaged to form the merged model.
-Model merging is based on the assumption that the models being merged lie in the same flat loss basin(Neyshabur et al., 2020), hence their interpolations also have low losses.
-Aside from its applications in multi-task learning, model merging can be naturally leveraged to reduce forgetting(Marczak et al., 2024).
-3 Analysis of Degradations in MCL
-Modality-incremental Continual Learning (MCL) is a special scenario of CL, where models incrementally learn on the data from new modalities.
-The first attempt of MCL is (Yu et al., 2024a).
-However, we point out that the MCL faces a more severe problem.
-In contrast to traditional CL scenarios, where the degradation comes solely from the forgetting of old knowledge, the degradation encountered in MCL comes from two aspects: forgetting and misalignment.
-Forgetting: the modality-agnostic components (the LLM backbone for MLLM) forget the knowledge of old modalities.
-Misalignment: the modality-agnostic components are misaligned with the modality-specific ones.
-When adapting to a new modality, the LLM backbone and the newly added modality encoder and connector are updated while the old ones are kept frozen.
-Therefore, the LLM backbone inevitably drifts away from the original multimodal feature space of old modality encoders (even the feature space of the new modality encoder, if its connector is not synchronously updated with LLM backbone).
-This process results in misalignment, further worsening the degradation.
-A potential exception to misalignment is when replay-based methods are applied since the old modality-specific components are also updated synchronously.
-However, empirical results in Section 5.5 suggest that replay-based methods still suffer from a certain degree of misalignment and it can be compensated by our proposed realigning stage.
-Due to the existence of misalignment, MCL problem requires special treatments compared to traditional CL problems.
-Figure 1: Pipeline of the proposed MERA. The procedures in gray boxes involve training. Refer to caption and Refer to caption represent the frozen and trainable modules, respectively.
+Recent advances (Panagopoulou et al., 2023; Chen et al., 2023a; Wu et al., 2024; Han et al., 2024; Zhan et al., 2024) in MLLM have extended LLMs to perceive multimodal inputs such as image, video, audio, point cloud, etc.
+Among all the MLLMs, the most influential one is LLaVA (Liu et al., 2024b, a), which utilizes a simple MLP connector to project visual information encoded by the pre-trained vision encoder into the language embedding space.
+Due to its simplicity and effectiveness, LLaVA-like architecture is widely adopted by a wide range of subsequent MLLMs (Lin et al., 2024b, a; Maaz et al., 2024; Wu et al., 2024; Chen et al., 2024) and accounts for the majority of current MLLMs.
+In this paper, we assume that the MLLM has a LLaVA-like architecture that includes modality-specific encoders and connectors, and a shared modality-agnostic LLM backbone.
+The rapid development of MLLMs demands high efficiency in their training process.
+It is efficient to reuse the existing MLLMs and extend them to more modalities.
+However, directly fine-tuning MLLMs on new modalities often results in significant performance degradation in previously learned modalities.
+In this work, we leverage the continual learning technique to tackle this problem.
+2.2 Traditional Continual Learning
+Continual Learning (CL) (Van de Ven and Tolias, 2019; Wang et al., 2024a) aims to continually acquire new knowledge with minor forgetting of previously learned knowledge.
+Existing CL methods mainly fall into the following four categories: Regularization-based methods (Kirkpatrick et al., 2017; Huszár, 2017; Schwarz et al., 2018) seek to protect the parameters that store important knowledge.
+However, storing an importance matrix during training requires extra memory with the same scale as the trainable parameters.
+Architecture-based methods (Yu et al., 2024a, b; Zadouri et al., 2024; Srinivasan et al., 2023) add task-specific parameters to the base model for each new task.
+This category requires modifications to the model architecture, harming its reusability.
+For many methods in this category, the model scale grows linearly as tasks increase, introducing extra memory overhead.
+Replay-based methods (Rolnick et al., 2019; Scialom et al., 2022; Wang et al., 2024b) leverage a small subset of historical data and replay it when learning on new data.
+This category requires access to partial data from previous tasks or distributions.
+However, this drawback is relatively minor in real applications since the replay data is often accessible.
+Merging-based methods (Wortsman et al., 2022; Marczak et al., 2024; Xiao et al., 2024; Zhu et al., 2024) edit models in parameter space to integrate the previously learned knowledge into the fine-tuned models by model merging (Yadav et al., 2024a; Yu et al., 2024c; Yang et al., 2024b, a).
+Although many traditional CL methods have been proposed, they are not specifically designed for Modality-incremental Continual Learning (MCL).
+In this work, we propose a simple MCL paradigm tailored for MLLMs.
+Table 1 summarizes the characteristics of different CL categories and our proposed method.
+2.3 Continual Learning for MLLMs
+Aside from traditional continual learning, there are methods tailored for MLLMs (Srivastava et al., 2024; Zeng et al., 2024; Maharana et al., 2024; Gao et al., 2024; Cao et al., 2024; Yu et al., 2024a).
+However, most of these works are specific to vision-language-only MLLMs, and incompatible with MCL where MLLMs can extend to arbitrarily more modalities.
+To the best of our knowledge, PathWeave (Yu et al., 2024a) is the most relevant work to MCL, and is so far the only work on MCL for MLLMs.
+It is an architecture-based method that uses an adapter-in-adapter mechanism to alleviate forgetting in previous modalities.
+In this work, we dive into the causes of degradation in MCL, uncovering that it suffers not only from forgetting but also from misalignment.
+To this end, we propose our two-stage MCL paradigm to address both forgetting and misalignment.
+Figure 1: Illustration of misalignment and the mechanism of our proposed realigning.
+$\phi_{i}$ is the feature distribution of the $i$-th modality.
+Regions in yellow represent the LLM’s expected distribution of the connector’s output.
+(a) and (b) are the states of the last learning stage and the ideal MCL after learning on a new modality.
+(c) shows the actual misalignment after learning on a new modality.
+(d) demonstrates the mechanism of our proposed realigning.
+3 Dual Causes of Degradation in MCL
+3.1 Preliminary
+We define the Modality-incremental Continual Learning (MCL) problem as follows.
+Given a sequence of $m$ modalities $\{M_{1},M_{2},\dots,M_{m}\}$ and their corresponding datasets $\{D_{1},D_{2},\dots,D_{m}\}$, MCL sequentially learns on each modality $M_{i}$ to obtain the model $\theta_{i}$.
+We denote the MLLM as $\theta=\{\theta^{Enc},\theta^{Conn},\theta^{LLM}\}$, where $\theta^{Enc},\theta^{Conn},\theta^{LLM}$ denote the modality encoders, the connectors, the LLM backbone, respectively.
+Notably, the $\theta^{Enc}$ and $\theta^{Conn}$ are modality-specific components and the $\theta^{LLM}$ is a modality-agnostic component.
+Further, we denote the feature distribution of the outputs of each modality connector $\theta_{i}^{Conn}$ as $\phi_{i}$.
+3.2 Forgetting and Misalignment in MCL
+MCL is a special scenario of CL, where models incrementally learn on new modalities.
+However, MCL faces a more severe problem.
+During continual learning, models would suffer from performance degradation in previously learned domains, tasks, or modalities.
+In traditional CL, degradation comes from the forgetting of previously learned knowledge.
+However, in MCL, it comes from two significant aspects: forgetting and misalignment.
+Forgetting: the modality-agnostic $\theta^{LLM}$ forgets the knowledge of old modalities.
+It is associated with various factors, including representation drift (Caccia et al., ), gradient interference (Wang et al., 2023), learning dynamics (Ren and Sutherland, 2024), distribution shift, etc.
+Misalignment: the modality-agnostic $\theta^{LLM}$ is misaligned with the modality-specific $\theta^{Enc}$.
+When incrementally learning on a new modality $M_{i}$, the $\theta^{LLM}$ is updated along with the modality connector of $M_{i}$, while other old connectors are kept frozen.
+Therefore, $\theta^{LLM}$ adapts to the new feature distribution $\phi_{i}$ and drifts away from the original multimodal distribution $\phi_{1}\cup\phi_{2}\cup\dots\cup\phi_{i-1}$ [^1] For our proposed method, the $\theta^{LLM}$ also drifts away from $\phi_{i}$ due to the use of model merging..
+Hence, there exists a misalignment in feature mapping between the old modality encoders and the $\theta^{LLM}$, leading to the breakdown of the "encoder-connector-LLM" collaboration chains for old modalities.
+The illustration of misalignment is sketched in Figure 1.
+Experiments in Section 6.1 also provide empirical evidence of the existence of misalignment.
+Above, we analyzed the dual causes of performance degradation in MCL. Due to the existence of misalignment, MCL problem requires special treatments in contrast to traditional CL problems.
 4 Method
-First, we define the Modality-incremental Continual Learning (MCL) problem as follows.
-Given a sequence of $m$ modalities $\{M_{1},M_{2},\dots,M_{m}\}$ and their corresponding datasets $\{D_{1},D_{2},\dots,D_{m}\}$, the model sequentially learns on each $D_{i}$ with the permission to access a $r\%$ size subset of previous data as replay data [^1] Different from the definition of replay data in the context of replay-based methods, where $R_{i}$ is sampled from $\{D_{1},D_{2},\dots,D_{i-1}\}$.
-In this paper, if not explained, $R_{i}$ for replay-based methods is sampled from $\{D_{1},D_{2},\dots,D_{i-1}\}$ while $R_{i}$ for our realigning stage is sampled from $\{D_{1},D_{2},\dots,D_{i}\}$.
-$R_{i}\leftarrow$ sample $r\%$ data from $\{D_{1},D_{2},\dots,D_{i}\}$.
-We denote the model after the $i$-th incremental training stage as $\theta_{i}=\{\theta_{i}^{agn},\theta_{i}^{spec}\}$, where $\theta_{i}^{agn},\theta_{i}^{spec}$ denote modality-agnostic and modality-specific components, respectively.
-For MLLM, the modality-agnostic component is the LLM backbone, and modality-specific components include encoders and connectors of each modality.
-To mitigate performance degradations in the previously learned modalities, we propose a two-stage MCL paradigm called "MErge then ReAlign" (MERA).
-In each stage of MCL, MERA executes the following two stages: merging and realigning, to address the forgetting and misalignment respectively.
-The overall pipeline of MERA is illustrated in Figure 1.
+To tackle the dual causes of performance degradation in MCL, we propose a two-stage MCL paradigm called "MErge then ReAlign" (MERA).
+In each stage of MCL, MERA executes the following two stages: merging and realigning, to address forgetting and misalignment respectively.
 4.1 Stage 1: Merging
-Motivation.
-Model merging has achieved great success in the field of multi-task learning for its superior capability of combining the strengths of multiple models into one.
+Model merging is efficient in integrating the previously learned knowledge into the fine-tuned models.
 Moreover, Yadav et al.
 (2024b) finds that model merging is more effective with larger models.
-Therefore, applying model merging to large-scale models, such as MLLMs, is inherently beneficial.
+Therefore, applying model merging to large-scale models such as MLLMs is inherently beneficial.
 Inspired by these, we introduce model merging to our MCL framework to mitigate forgetting.
-However, model merging methods generally can not be directly applied to continual learning.
-In this work, we only focus on the simplest model merging method, i.e., weight averaging, and revise it into a CL form to provide a basic framework.
-To adapt weight averaging to CL, we associate its merging coefficients with the progress of CL stages, forming a Cumulative Moving Average (CMA) merging.
-At the $i$-th training stage, the CMA merged model is calculated by:
-$CMA(\theta_{i-1},\theta_{i};i)=\frac{i-1}{i}\theta_{i-1}+\frac{1}{i}\theta_{i}$
-To further adapt to the MCL framework, we only merge the modality-agnostic components (the LLM backbone) and ensemble the modality-specific components (encoders and connectors of each modality):
-[…]
-$\theta_{i-1}^{spec},\theta_{i}^{spec}\}$
-To perform MCL, when starting from $\theta_{i-1}$ to incrementally learn a new modality $M_{i}$, firstly we directly train the model on $D_{i}$ to get $\theta_{i,vanilla}$.
-Note that the training process here is the same as the standard way of training an MLLM on a single modality, except that the initial weights of the model are inherited from $\theta_{i-1}$.
-Then secondly, we apply CMA merging to $\theta_{i,vanilla}$ and $\theta_{i-1}$ to derive the merged model $\theta_{i,merged}=CMA^{*}(\theta_{i-1},\theta_{i,vanilla};i)$ with integrated knowledge of both new modality and old modalities.
-Discussion.
-Different from other CL methods, the LLM backbone is asynchronously updated with the new connector in the merging stage.
-To be specific, in the standard MLLM training step, they are synchronously updated indeed, however, in the CMA merging step, only the LLM backbone is updated, resulting in an additional misalignment between them [^2] This is the reason we sample $R_{i}$ from $\{D_{1},D_{2},\dots,D_{i}\}$ rather than $\{D_{1},D_{2},\dots,D_{i-1}\}$ in our realigning stage..
+The first step before model merging is to perform the standard MLLM training, which often encompasses a pre-training and a fine-tuning phase.
+After the standard training step, we get the vanilla model $\theta_{i,vanilla}$ that inevitably suffers from forgetting the knowledge of previous modalities.
+The second step is to perform model merging to mitigate forgetting.
+In this work, we only focus on the simplest model merging method, i.e., weight averaging, aiming only to provide a basic framework.
+To adapt weight averaging to MCL, we associate its merging coefficients with the progress of MCL stages.
+At the $i$-th training stage, the merged model is calculated by:
+$\theta_{i,merged}=\frac{i-1}{i}\theta_{i-1}+\frac{1}{i}\theta_{i,vanilla}$
+Notably, we only merge the modality-agnostic component $\theta^{LLM}$ and ensemble the modality-specific components $\theta^{Enc}$ and $\theta^{Conn}$.
+After merging, we obtain the $\theta_{i,merged}$, whose knowledge of previous modalities is enhanced.
 4.2 Stage 2: Realigning
-To address the misalignment issue, we simply leverage a small replay dataset $R_{i}\leftarrow$ sample $r\%$ data from $\{D_{1},D_{2},\dots,D_{i}\}$ to further fine-tune all the connectors of $\theta_{i,merged}$ to obtain the final $\theta_{i}$.
-This process realigns the encoders of each modality with the LLM backbone.
-The training objective for realigning is unchanged from the original MLLM training objective, i.e., the auto-regressive loss.
-The realigning stage is similar to replay-based CL methods(Scialom et al., 2022; Wang et al., 2024) in form as they both leverage a replay dataset, however, they are essentially different.
+To address the misalignment issue, we propose a lightweight realigning stage to update the multimodal distribution $\phi_{1}\cup\phi_{2}\cup\dots\cup\phi_{i}$ to realign the "encoder-connector-LLM" chains for all the modalities.
+The realigning stage simply leverages a small replay dataset [^2] It is different from replay data in replay-based continual learning methods, where their $R_{i}$ is sampled from $\{D_{1},D_{2},\dots,D_{i-1}\}$.
+$R_{i}\leftarrow$ sample rr% data from $\{D_{1},D_{2},\dots,D_{i}\}$ to further fine-tune all the connectors of $\theta_{i,merged}$.
+This realigning process is formulated as:
+$\min\limits_{\theta_{i,merged}^{Conn}}\mathbb{E}_{x\sim R_{i}}\mathcal{L}(\theta_{i,merged},x)$
+where the $\mathcal{L}$ is the auto-regressive loss, unchanged from the original MLLM training loss.
+By fine-tuning the lightweight $\theta^{Conn}$ with only a small replay dataset, it efficiently realigns the $\theta^{Enc}$ with $\theta^{LLM}$.
+After realigning, we can obtain the final model $\theta_{i}$.
+Figure 2: Pipeline of the proposed MERA. The procedures in gray boxes involve training. Refer to caption and Refer to caption represent the frozen and trainable modules, respectively.
+Differences between realigning and replay-based CL methods.
+The realigning stage resembles replay-based CL methods (Scialom et al., 2022; Wang et al., 2024b) in form as they both leverage replay data, however, they are essentially different.
 First, replay methods train on the joint dataset of $D_{i}$ and $R_{i}$, while our realigning stage trains solely on $R_{i}$.
-Second, the purpose of replay is to review the previous knowledge, while the motivation of realigning stage is to align modality spaces.
-Third, replay methods update all the parameters to consolidate old knowledge, in contrast, keeping the LLM backbone frozen is crucial in the realigning stage, otherwise the LLM would overfit on replay data.
+Second, replay methods update both the $\theta^{LLM}$ and $\theta^{Conn}$, while the realigning stage only efficiently updates the $\theta^{Conn}$ to prevent the knowledge inside $\theta^{LLM}$ from being overwritten again.
+4.3 Overall Pipeline
+The overall pipeline of MERA is illustrated in Figure 2.
+In each stage of MCL, MERA goes through two stages.
+The first stage is merging, where we fine-tune $\theta_{i-1}$ on the incoming modality and merge the fine-tuned model $\theta_{i-1,vanilla}$ with the historical model $\theta_{i-1}$ in order to alleviate forgetting.
+The second stage is realigning, where we leverage a small set of replay data $R_{i}$ to efficiently fine-tune the lightweight modality connectors $\theta^{Conn}$ to realign the $\theta^{Enc}$ with $\theta^{LLM}$.
 5 Experiments
@@
 Reverse Order: point cloud $\rightarrow$ audio $\rightarrow$ video $\rightarrow$ image.
-On top of this, the adopted datasets, metrics, models, and baselines are listed as follows.
+On top of this, the adopted datasets, metrics, models, and baselines are detailed as follows.
 Datasets.
@@
 For point cloud modality, we use a subset of Cap3D (Luo et al., 2024) and a subset of Cap3D-QA (Panagopoulou et al., 2023).
-More details of these datasets are in Appendix A.
+More details of these datasets are in Appendix B.
 Evaluation Metrics.
-First, we leverage Relative Gain(Scialom et al., 2022; Wang et al., 2024) as a normalized metric across different tasks.
-We train expert MLLMs individually on each single modality $M_{j}$ and test with their respective holdout data, taking their scores on the $k$-th dataset $D_{j,k}$ as upper bound $S_{j,k}^{sup}$.
+First, we leverage Relative Gain (Scialom et al., 2022; Wang et al., 2024b) as a normalized metric across different tasks.
+We naively train (without using any continual learning methods) expert MLLMs individually on each single modality $M_{j}$ and test with their respective holdout data, taking their scores on the $k$-th dataset $D_{j,k}$ as upper bound $S_{j,k}^{sup}$.
 In the incremental stage $i$, the Relative Gain of modality $M_{i}$ with its dataset $D_{j}=\{D_{j,k}\}_{k=1}^{K}$ is calculated by:
@@
 Here, we utilize CIDEr score (Vedantam et al., 2015) and prediction accuracy (Acc) for Cap and QA tasks respectively to calculate $S_{j,k}^{sup}$ and $S_{j,k}^{i}$.
-To evaluate the performance degradations of the previously learned modalities, we calculate the Backward Relative Gain in the stage $i$ as:
+To evaluate the performance degradation of the previously learned modalities, we calculate the Backward Relative Gain in the stage $i$ as:
 $\mathrm{Bw\ Relative\ Gain}^{i}=\frac{1}{i-1}\sum_{j=1}^{i-1}\mathrm{Relative\ Gain}_{j}^{i}$
@@
 $\mathrm{Fw\ Relative\ Gain}^{i}=\mathrm{Relative\ Gain}_{i}^{i}$
+Figure 3: Progressive Backward Relative Gain in modality-incremental continual learning.
+For each stage $i$, we plot the average score of the corresponding Backward Relative Gain with two different training orders.
+We set Backward Relative Gain to 100% for the 11st stage, denoting the initial performance without degradation.
+Exceptionally, the initial Backward Relative Gain of EProj is not 100% since it only tunes the modality-specific components, causing an initial performance degradation.
 Model and Training Details.
 We leverage the mainstream MLLM architecture, i.e., LLaVA-like architecture with the Llama-3-8B-Instruct (Dubey et al., 2024) as its LLM backbone.
-The selections of modality encoders and connectors are detailed in Appendix B.
+The selections of modality encoders and connectors are detailed in Appendix C.1.
 Trainings that involve updating the LLM backbone utilize LoRA (Hu et al., 2022) for parameter-efficient fine-tuning.
-The training process in our merging stage is the same as the regular MLLM training, i.e., in the first step, only the connector is updated with Cap datasets, then in the second step, the connector and the LLM backbone are updated with all the task-related datasets (the combination of Cap and QA datasets in our case).
+The training process in our merging stage is the same as the standard MLLM training, i.e., in the first step, only the connector is updated with Cap datasets, then in the second step, the connector and the LLM backbone are updated with all the task-related datasets (the combination of Cap and QA datasets in our case).
 In our realigning stage, the replay datasets are randomly sampled from the joint datasets of Cap and QA tasks.
-For each training process, the hyperparameters are listed in Appendix B.
-Figure 2: Progressive Backward Relative Gain in modality-incremental continual learning.
-For each stage $i$, we plot the average score of corresponding Backward Relative Gain with two different training orders.
-We set Backward Relative Gain to 100% for the $1$-st stage, denoting the initial performance without degradation.
+For each training process, the hyperparameters are listed in Appendix C.1.
 Baselines.
-In our experiments, we compare our MERA with non-CL fine-tuning and the representative methods of each CL category: Fine-Tuning: directly train MLLMs sequentially on each modality without applying any CL method.
+In our experiments, we compare our MERA with non-CL fine-tuning, as well as the representative CL and MCL methods: Fine-Tuning: directly train MLLMs sequentially on each modality without applying any CL method.
 Replay: the vanilla replay-based CL method.
-During training on a new task, the model is updated with both samples from the current task and a set of randomly sampled replay data from previous tasks, to review knowledge of earlier tasks while learning the new one.
+During training on a new task, the model is updated with both samples from the current task and a set of randomly sampled replay data from previous tasks.
 EWC (Kirkpatrick et al., 2017): the most representative regularization-based CL method.
@@
 It uses the Fisher information matrix to measure the importance of each weight.
-PathWeave(Yu et al., 2024a): an architecture-based CL method, also the first MCL method for MLLMs.
+EProj (He et al., 2023): tuning only the modality-specific components to prevent forgetting.
+PathWeave (Yu et al., 2024a): an architecture-based CL method, also the state-of-the-art MCL method for MLLMs.
 PathWeave uses an adapter-in-adapter mechanism to memorize and extract knowledge from historical modalities to enhance the learning of the current modality.
 PathWeave is originally built on X-InstructBLIP (Panagopoulou et al., 2023).
-For a fair comparison, we implement PathWeave for our adopted MLLM architecture [^3] In their original settings, PathWeave removes the newly added modules when testing the former modalities.
-However, this results in the inability to perform cross-modality tasks, which are common in real applications.
-Therefore, we do not remove them for a fair comparison..
-Implementation details of each baseline method are in Appendix D.
+For a fair comparison, we implement PathWeave for our adopted MLLM architecture.
+Implementation details of each baseline method are in Appendix C.3.
+Figure 4: Progressive Forward Relative Gain in modality-incremental continual learning.
+For each stage $i$, we plot the average score of the corresponding Forward Relative Gain with two different training orders.
+We set Forward Relative Gain to 100% for the 11st stage, denoting the initial lossless plasticity.
+Exceptionally, the initial Forward Relative Gain of EProj is not 100% since it only tunes the modality-specific components, causing an initial loss of plasticity.
+5.2 Main Results
+We conduct experiments under our MCL setting with both sequential and reverse orders.
+For Replay and our MERA, results using rr%, $r=\{1,10\}$ replay data are reported, denoted by Replay (rr%) and MERA (rr%) respectively.
+Evaluating Degradation.
+The progressive Backward Relative Gains averaged from different training orders are plotted in Figure 3.
+It is observed that our MERA demonstrates an impressive capability of mitigating performance degradation with consistent and promising Backward Relative Gains.
+When extending to all four modalities, MERA ($10\%$) holds up to a 99.84% Backward Relative Gain, indicating that MERA can achieve nearly lossless MCL performance, with at least 12.09% absolute improvements of Backward Relative Gain compared with other baselines.
+Notably, when only leveraging $1\%$ replay data, MERA (1%) can still achieve at least 7.66% absolute improvements over other baselines.
+Further, we calculated the mean and standard deviation of Backward Relative Gains in all training stages for each method, in different training orders.
+Table 2 shows that our MERA (10%) achieves the highest mean in both training orders, indicating its superior performance.
+It also achieves the lowest and second-lowest standard deviation in sequential and reverse orders respectively, indicating its high stability.
+Notably, in sequential order, MERA (10%) performs even better than lossless MCL, with an over 100% average Backward Relative Gain, also at least 11.39% absolute improvements of average Backward Relative Gain over other baselines.
+In reverse order, MERA (10%) also achieves at least 13.33% absolute improvements.
+When with only 1% replay data, MERA (1%) still achieves at least 8.29% and 4.33% absolute improvements in sequential and reverse orders respectively.
 Method Sequential Reverse
@@
 EWC 74.93 17.14 68.01 21.54
+EProj 89.61 2.63 70.07 11.86
 PathWeave 86.85 12.17 80.09 13.31
@@
 The best results are in bold, while the second-best are underlined.
+Method Sequential Reverse
+Fine-Tuning 38.50 36.14
+Replay (1%) 30.61 33.02
+Replay (10%) 56.68 32.50
+EWC 25.79 24.94
+EProj 55.61 55.61
+PathWeave 36.32 46.39
+MERA (1%) 72.13 61.22
+MERA (10%) 72.70 54.04
+Table 3: Accuracies on MCUB-4 benchmark. Results are reported on different training orders and are measured after continually learning on all four modalities.
+Evaluating Degradation on Complex Multimodal Tasks.
+To evaluate MERA’s performance in more complex multimodal tasks that involve multiple modalities at a time, we further conduct experiments on the MCUB-4 (Chen et al., 2024) benchmark that requires the model to simultaneously infer on all four modalities, i.e., image, video, audio, and pointcloud.
+Table 3 shows the accuracies of each method on MCUB-4 benchmark.
+It is observed that our proposed MERA also significantly excels in more complex multimodal tasks that involve multiple modalities at a time.
+Evaluating Plasticity.
+Aside from alleviating performance degradation, the capability to adapt to new knowledge, i.e., plasticity, is also an important aspect.
+We use the Forward Relative Gain as the metric.
+The progressive Forward Relative Gains averaged from different training orders are plotted in Figure 4.
+It is observed that the most elastic CL methods are EWC and PathWeave, while our MERA (10%) demonstrates comparable plasticity.
+5.3 Ablation Study
+We conduct ablation studies to investigate the effectiveness of each stage of MERA.
+Results are shown in Table 4.
+Firstly, from Table 4 and Table 2, it is observed that the realigning stage that addresses the misalignment issue can already beat many other baselines designed to tackle the forgetting issue, achieving the second-best average Backward Relative Gain among baselines in both sequential and reverse orders.
+Secondly, combining both merging and realigning stages, MERA further narrows the gap between the incrementally learned models and the individually trained experts on each modality, even, surpassing the individually trained experts in sequential training order with over 100% Backward Relative Gain.
+Method Sequential Reverse
+Mean Std Mean Std
+Fine-Tuning 59.76 27.23 48.96 35.70
++Merging 90.29 7.42 70.00 30.87
++Realigning 87.92 12.41 71.90 24.52
+MERA 101.00 3.90 93.42 6.25
+Table 4: Ablation study of different components in MERA. The realigning stage uses 10% replay data.
+6 Discussions
+6.1 Is Misalignment Common in MCL?
+Since the realigning stage achieves great success on top of our proposed merging stage, we further ask another question: does realigning benefit other CL methods, or in other words, is misalignment a common phenomenon in MCL?
+To examine this, we perform the realigning stage at the end of every training stage for different CL or non-CL methods to observe whether there are performance improvements [^3] We do not examine this for EProj.
+EProj is self-evidently misalignment-free, since it freezes the LLM backbone..
+Table 5 shows that the additional realigning stage brings substantial performance improvements and increased stability for different CL or non-CL methods.
+Based on this observation, we can conclude that misalignment is a common phenomenon in MCL, and can be alleviated by our proposed realigning stage.
+Method Sequential Reverse
+Mean Std Mean Std
+Fine-Tuning 59.76 27.23 48.96 35.70
++Realigning +28.16 -14.83 +22.93 -11.17
+Replay (1%) 66.09 25.86 43.95 39.03
++Realigning +20.64 -9.75 +24.35 -8.36
+Replay (10%) 77.52 16.32 51.90 36.34
++Realigning +14.21 -6.67 +23.71 -11.07
+EWC 74.93 17.14 68.01 21.54
++Realigning +19.54 -6.87 +23.02 -13.25
+PathWeave 86.85 12.17 80.09 13.31
++Realigning +6.22 -0.28 +2.40 -1.53
+Merging 90.29 7.42 70.00 30.87
++Realigning +10.71 -3.52 +23.42 -24.63
+Table 5: Applying realigning to different CL methods can further improve their Backward Relative Gain and stability. The realigning stage uses 10% replay data. Improvements are colored in green.
+6.2 Positive Backward Transfer and Positive Forward Transfer
+From Figure 3 and its raw data shown in Appendix D, we observe a faint phenomenon of Positive Backward Transfer (Lin et al., 2022) that learning new knowledge improves the performance on previously learned tasks.
+For most methods, the Backward Relative Gain comes to a low level when incrementally learning the second modality, but starts to stabilize and even increase when incrementally learning more modalities.
+From Figure 4, we observe a strong phenomenon of Positive Forward Transfer (Ke et al., 2021) that the knowledge acquired from earlier tasks improves the learning efficiency of new tasks.
+The Positive Forward Transfer exists before the $4$-th incremental stage, when employing EWC, PathWeave, and MERA.
+This phenomenon is also reported in other MCL literature (Yu et al., 2024a).
+In contrast to Positive Forward Transfer, there is a gradual loss of plasticity (Dohare et al., 2024, 2023) as the model attempts to retain more knowledge.
+This explains the decreases in Forward Relative Gain across different CL methods in the $4$-th stage, as the loss of plasticity comes to a dominant position.
+7 Future Work
+This work is one of the early attempts of MCL for MLLMs, focusing on the dual causes of its degradation.
+In addition to this, observation of Positive Backward Transfer and Positive Forward Transfer in Section 6.2 may imply the complex cross-modal interaction in multimodal learning, urging for future research on the mechanisms of modality interaction in the context of MCL.
+8 Conclusion
+In this paper, we first revisit MCL and uncover the dual causes of its degradation, i.e., forgetting and misalignment.
+Next, to address both forgetting and misalignment, we propose MERA, a simple yet effective MCL paradigm.
+Extensive experiments demonstrate that MERA significantly outperforms the state-of-the-art methods and even achieves nearly lossless MCL performance.
+Our findings underscore the misalignment issue in MCL.
+More broadly, our work showcases how to adjust different components of MLLMs during continual learning.
+Further, we observe signs of complex cross-modal interaction in MCL, providing a direction for future work.
+9 Limitations
+This work is restricted in the following aspects.
+First, our experiments are limited to four commonly used modalities due to the lack of resources for other less-studied modalities.
+Second, we limit this work to LLaVA-like architecture as it covers the majority of MLLMs.
+Third, this work is limited to any-to-text MLLMs while there is now a trend of exploring any-to-any MLLMs.
+However, the main idea of MERA is generic to them.
+Fourth, our work only explores the simplest model merging method in the context of MCL, aiming to provide a universal framework, leaving other model merging methods for MCL for future work.
+10 Acknowledgments
+This work was supported by the National Natural Science Foundation of China (No.62372139), the National Natural Science Foundation of China (2024A1515030024), Research Projects of Shenzhen (JCYJ20220818102414030), and Key Laboratory of Guangdong Province (2022B1212010005).
+Appendix
+Appendix A Further Analysis and Discussions
+A.1 Deeper Discussions on the Distinction Between Forgetting and Misalignment
+Theoretically, forgetting is associated with various factors such as representation drift (Caccia et al., ), gradient interference (Wang et al., 2023), learning dynamics (Ren and Sutherland, 2024), distribution shift (which causes misalignment), making it a comprehensive issue.
+For example, Schwettmann et al.
+(2023) finds that there exist multimodal neurons in the LLM backbone, each corresponds to certain domain concepts, and when breaking some of these neurons, the model’s prediction is largely affected.
+However, addressing the misalignment issue by our proposed realigning will possibly not recover the model’s performance since the connectors do not encode domain-specific concepts (Verma et al., 2024).
+This example showcases that the forgetting issue is not limited to misalignment.
+Empirically, our experiments in Table 5 have shown that, a) addressing only the misalignment issue, i.e., fine-tuning with realigning, does not outperform some methods that address the overall forgetting, e.g., PathWeave/merging without realigning, b) methods that address the overall forgetting significantly complement the realigning alone (fine-tuning with realigning).
+In terms of methodology, forgetting is a comprehensive issue that can be tackled mainly by heuristic methods such as EWC or replay, while misalignment is caused solely by distribution shift that can be directly addressed by realigning the LLM and modality-specific components.
+In terms of methodology, forgetting is a comprehensive issue that can be tackled mainly by heuristic methods such as EWC or replay, while misalignment is caused solely by distribution shift that can be directly addressed by realigning the LLM and modality-specific components.
+A.2 Misalignment Also Exists In Replay-Based Methods
+Theoretically, a potential exception to misalignment is when replay-based methods are applied, since the $\theta^{LLM}$ is trained on the joint distribution of all $\{M_{1},M_{2},\dots,M_{i}\}$ modalities.
+However, Table 5 suggests that replay-based methods still suffer from misalignment and can be compensated for by our proposed realigning stage.
+We conjecture that it is due to the imbalanced distribution of its training data.
+Replay-based methods train the model on the joint dataset of its replay data and $D_{i}$, where the scale of replay data from each previous modality is insignificant to the scale of $D_{i}$.
+Therefore, it still suffers from a certain degree of distribution shift during its training process.
+A.3 Easily Adapt MERA to Other MLLM Architectures
+Although we limit our work to LLaVA-like architecture, our method can be easily adapted to other MLLM architectures. We present several naive ways to adapt MERA to some other MLLM architectures.
+  • For connector-free MLLMs, the last few layers of their encoders can be treated as connectors so that our method can be directly applied.
+  • For MLLMs that use a uni-connector for all modalities, we can treat them as the connector-free MLLMs.
+  • For encoder-free MLLMs, we can realign the raw multimodal input distributions instead of multimodal feature distributions with the LLM backbone by fine-tuning the embedding layers.
+A.4 Efficiency Comparisons
+We compare the efficiency of different baselines and our MERA, as shown in Table 6.
+It is observed that our MERA can achieve optimal results except that MERA (1%) and MERA (10%) introduce 2% and 15% extra training time-consuming respectively.
+However, we believe its trade-off between training time-consuming and performance is worthwhile, considering the impressive performance of MERA.
+In our experiments, EWC and PathWeave introduce marginal extra training memory overhead, as we employ parameter-efficient fine-tuning.
+However, for larger LoRA ranks or even full model fine-tuning, their extra training memory consumptions would be substantial, as they necessitate storing additional parameters whose sizes increase linearly with the trainable parameters.
 Method Training Inference
@@
 Mean Std Mean Std
-Fine-Tuning 59.76 27.23 48.96 35.70
-+Merging 90.29 7.42 70.00 30.87
-+Realigning 87.92 12.41 71.90 24.52
-MERA 101.00 3.90 93.42 6.25
-Table 4: Ablation study of different components in MERA.
-The realigning stage uses 10% replay data.
-The mean and standard deviation of Backward Relative Gains in all the training stages are reported on different training orders.
-The best results are in bold.
-5.2 Main Results
-We conduct experiments under our MCL setting with both sequential and reverse orders.
-For Replay and our MERA, results using $r\%,r=\{1,10\}$ replay data are reported, denoted by Replay ($r\%$) and MERA ($r\%$) respectively.
-The progressive Backward Relative Gains averaged from different training orders are plotted in Figure 2.
-It is observed that our MERA demonstrates an impressive capability of mitigating performance degradation with consistent and promising Backward Relative Gains.
-When extending to all the four modalities, MERA ($10\%$) holds up to a 99.84% Backward Relative Gain, indicating that MERA can achieve a nearly lossless MCL performance, with at least 20.37% absolute improvements of Backward Relative Gain compared with other baselines.
-Notably, when only leveraging $1\%$ replay data, MERA (1%) can still achieve at least 15.94% absolute improvements over other baselines.
-Further, we calculated the mean and standard deviation of Backward Relative Gains in all training stages for each method, in different training orders.
-Table 2 shows that our MERA (10%) achieves the highest mean and lowest standard deviation in both training orders, indicating its superior performance and high stability.
-Notably, in sequential order, MERA (10%) performs even better than lossless MCL, with an over 100% average Backward Relative Gain, also at least 14.15% absolute improvements of average Backward Relative Gain over other baselines.
-In reverse order, MERA (10%) also achieves at least 13.33% absolute improvements.
-When with only 1% replay data, MERA (1%) still achieves at least 11.05% and 4.33% absolute improvements in sequential and reverse orders respectively.
-From Figure 2 and its raw data shown in Appendix E, we also observe a faint phenomenon of Positive Backward Transfer(Lin et al., 2022) that learning new knowledge improves the performance on previously learned tasks.
-For most CL methods and non-CL fine-tuning, the Backward Relative Gain comes to a low level when incrementally learning the first modality (corresponding to the $2$-nd training stage), but starts to stabilize or even increase when incrementally learning more modalities.
-5.3 Efficiency Comparisons
-We further compare the efficiency of different baselines and our MERA, as shown in Table 3.
-It is observed that our MERA can achieve optimal results except that MERA (1%) and MERA (10%) introduce 2% and 15% extra training time-consuming respectively.
-However, we believe its trade-off between training time-consuming and performance is worthwhile, considering the impressive performance of MERA.
-In our experiments, EWC and PathWeave introduce marginal extra training memory overhead, as we employ parameter-efficient fine-tuning.
-However, for larger LoRA ranks or even full model fine-tuning, their extra training memory consumptions would be substantial, as they necessitate storing additional parameters whose sizes increase linearly with the trainable parameters.
-5.4 Ablation Study
-We conduct ablation studies to investigate the effectiveness of each component in MERA.
-Results are shown in Table 4.
-Firstly, from Table 4 and Table 2, it is observed that the merging stage alone can already beat many other baselines, achieving the best and second-best performances among baselines in sequential and reverse orders respectively.
-Secondly, the realigning stage alone can also achieve performance improvements by addressing the misalignment issue.
-Combining both the merging and realigning stages, MERA further narrows the gap between the incrementally learned models and the individually trained experts on each modality, even, surpassing the individually trained experts in sequential training order with over 100% Backward Relative Gain.
-Method Sequential Reverse
-Mean Std Mean Std
-Fine-Tuning 59.76 27.23 48.96 35.70
-+Realigning +28.16 -14.83 +22.93 -11.17
-Replay (1%) 66.09 25.86 43.95 39.03
-+Realigning +20.64 -9.75 +24.35 -8.36
-Replay (10%) 77.52 16.32 51.90 36.34
-+Realigning +14.21 -6.67 +23.71 -11.07
-EWC 74.93 17.14 68.01 21.54
-+Realigning +19.54 -6.87 +23.02 -13.25
-PathWeave 86.85 12.17 80.09 13.31
-+Realigning +6.22 -0.28 +2.40 -1.53
-Merging 90.29 7.42 70.00 30.87
-+Realigning +10.71 -3.52 +23.42 -24.63
-Table 5: Applying realigning to different CL or non-CL methods can further improve their Backward Relative Gain and stability.
-The realigning stage uses 10% replay data.
-The increased mean indicates better performance, while the decreased standard deviation indicates better stability.
-Improvements are colored in green.
-Figure 3: Progressive Forward Relative Gain in modality-incremental continual learning.
-For each stage $i$, we plot the average score of corresponding Forward Relative Gain with two different training orders.
-We set Forward Relative Gain to 100% for the $1$-st stage, denoting the initial lossless plasticity.
-5.5 Misalignment Is Common in MCL
-Since the realigning stage achieves great success on top of our proposed merging stage, we further ask another question: does realigning benefit other CL methods, or is misalignment a common phenomenon in MCL?
-To examine this, we perform the realigning stage at the end of every training stage for different CL or non-CL methods to observe whether there are performance improvements.
-Table 5 shows that the additional realigning stage brings substantial performance improvements and increased stability for different CL or non-CL methods.
-Based on this observation, we can conclude that misalignment is a common phenomenon in MCL, and can be compensated by our proposed realigning stage.
-5.6 Plasticity Analysis
-Aside from alleviating performance degradation, the capability to adapt to new knowledge, i.e., plasticity, is also an important aspect.
-We use Forward Relative Gain as the metric to evaluate it.
-The progressive Forward Relative Gains averaged from different training orders are plotted in Figure 3.
-It is observed that the most elastic CL methods are EWC and PathWeave, while our MERA (10%) demonstrates comparable plasticity.
-Notably, from Figure 3, we observe a strong phenomenon of Positive Forward Transfer(Ke et al., 2021) that the knowledge or skills acquired from earlier tasks improve the learning efficiency of new tasks.
-The Positive Forward Transfer emerges before the $4$-th training stage, on EWC, PathWeave, and MERA.
-This phenomenon is also reported by other MCL literature(Yu et al., 2024a).
-In contrast to Positive Forward Transfer, there is a gradual plasticity loss(Dohare et al., 2024, 2023) as the model attempts to retain more knowledge.
-This explains the decreases in Forward Relative Gain across different CL methods in the $4$-th stage, as the plasticity loss comes to a dominant position.
-6 Future Work
-This work is one of the early attempts of MCL, mainly focusing on designing effective methods and addressing the misalignment issue.
-However, our extensive experiments suggest that there are many other interesting aspects of MCL that warrant attention.
-In Section 5.2 and Section 5.6, we observe the phenomenons of Positive Backward Transfer and Positive Forward Transfer, respectively.
-This could be related to the complex cross-modal interaction in multimodal learning, urging for further research on the mechanisms of modality interaction and methods that boost a positive modality interaction, in the context of MCL.
-7 Conclusion
-In this paper, we mainly focus on designing effective methods for Modality-incremental Continual Learning (MCL).
-First, we revisit MCL and investigate a more severe issue it faces in contrast to traditional continual learning that its performance degradation comes not only from forgetting but also from misalignment.
-To address both the forgetting and misalignment, we propose MERA, a simple yet effective MCL paradigm.
-Extensive experiments demonstrate that MERA significantly outperforms the baselines, and even achieves a nearly lossless MCL performance.
-Further, we observe a sign of complex cross-modal interaction in MCL, providing a direction for future work.
-Limitations
-Our work is restricted in the following aspects.
-First, our experiments are limited to four commonly used modalities due to the lack of resources for other less-studied modalities.
-Second, our experiments are limited to two types of tasks for each modality due to the lack of general-purpose and diverse instruction datasets and their corresponding benchmarks for some modalities, i.e., audio and point cloud.
-Third, our method is limited to any-to-text MLLMs while there is now a trend of exploring any-to-any MLLMs.
-However, the main idea of MERA is generic to any-to-any MLLMs since their architecture topologies are similar.
-Fourth, our work only explores the simplest model merging method in the context of MCL, aiming to provide a universal framework, leaving the adaptation of other model merging methods to MCL for future work.
-Appendix
-Appendix A Dataset Details
-Table 6 details the statistics of each dataset. Some datasets are filtered from their original ones:
-  • MSCOCO-2014(Lin et al., 2014): Each image has multiple captions, we only use its first caption to form the training set.
-  • Clotho-AQA(Lipping et al., 2022): Each sample is annotated with a confidence level, we only use the samples whose confidence levels are "yes" to form the training set and test set.
-  • Cap3D(Luo et al., 2024): Since the original dataset is huge in scale, we filter out the samples whose caption is longer than 100 letters. Then, we randomly sample a 50K subset as the training set.
-  • Cap3D-QA(Panagopoulou et al., 2023): Since the original dataset is huge in scale, we randomly sample a 30K subset as the training set.
-For each dataset, we use a randomly sampled 1K subset of its holdout test set as the final test set, except for the MSVD(Chen and Dolan, 2011), since the size of its original test set is less than 1K.
-Appendix B Implementation Details
-We build our experimental codebase on top of LLaVA(Liu et al., 2024b, a) and NExT-GPT(Wu et al., 2024). We detail the modality-specific components of each modality as follows:
-  • Image: We use CLIP-ViT-L-336px(Openai, 2021) as the pre-trained image encoder, a randomly initialized MLP as the connector.
-  • Video: We use CLIP-ViT-L-336px(Openai, 2021) as the pre-trained video encoder, a randomly initialized MLP as the connector.
-We uniformly sample 4 frames from a video as input frames.
-Then each frame is encoded by the video encoder separately.
-The output feature frames are downsampled by 2x using bilinear pooling before sending into the connector to improve efficiency.
-[…]
-  • Point Cloud: We use Point-BERT-v1.2Xu et al. (2024) as the pre-trained point cloud encoder, a randomly initialized MLP as the connector.
-We set the hyperparameters mainly following previous worksLiu et al., 2024b, a, as listed in Table 7.
-For training that involves updating the LLM backbone, we utilize parameter-efficient fine-tuning with LoRA(Hu et al., 2022) applied across all linear modules within the LLM, setting the LoRA rank to 128 and the alpha parameter to 128.
-All the experiments are conducted on a single NVIDIA RTX A6000 48G with FP16.
+MERA 97.90 6.02 84.42 12.93
+MERA
+w/ 10% noise
+96.87 4.97 83.56 13.48
+MERA
+w/ 50% noise
+94.10 5.67 79.79 15.54
+Table 7: Evaluations of MERA’s robustness to noisy replay data. The realigning stage uses 10% replay data.
+A.5 Robustness to the Quality of Replay Data
+Since the realigning stage of MERA relies on a small replay dataset, it is beneficial to understand how robust MERA is to the quality of replay data.
+To evaluate its robustness, we manually corrupt $p\%$ samples in the replay dataset by mispairing their text-modality pair, and test MERA’s performance.
+Results in Table 7 suggest that 10% noisy samples does not significantly corrupt MERA’s performance, and it still shows decent performance even under the extreme condition of 50% noisy samples.
+This suggests that our MERA is robust to low-quality replay data.
+A.6 Why is the Performance Degradation More Severe in Reverse Training Order
+In Table 2, it is observed that the performance degradation is more severe in reverse training order than in sequential training order.
+This phenomenon occurs across different methods.
+We conjecture that training in sequential order is a sort of curriculum learning, while the reverse order corresponds to reversed curriculum learning.
+In our experiments, the sequential order is determined based on the prevalence of different modalities, and more prevalent modalities might be easier to learn.
+For example, the image modality is easier to understand than video, and the point cloud might be the most difficult modality to learn.
+Wang et al.
+(2024b) also observed that the performance degradation is less severe when continually learning in a curriculum learning order than in a reversed curriculum learning order.
+We conjecture that learning easy knowledge leads to the forgetting of harder ones, but learning hard knowledge might even consolidate the easier ones.
 #Training Set #Test Set License
@@
 Table 8: Statistics of the datasets. Datasets marked with * are filtered from their original ones.
-Hyperparameters Std-Stage 1 Std-Stage 2 Realigning
+Hyperparameters Pre-Training Fine-Tuning Realigning
 Trainable Components Connectors LLM and Connectors Connectors
@@
 Epoch 1
-Table 7: Hyperparameters for each training stage. Std-stage 1 and Std-stage 2 refer to the stage 1 and 2 of the standard MLLM training process.
-Appendix C Evaluation Details
-For the calculation of CIDEr scores(Vedantam et al., 2015), we utilized an open-sourced library CaptionMetrics(wangleihitcs, 2019).
-For the calculation of prediction accuracy, we leverage a GPT-based open-ended QA evaluation with GPT-4o mini as the judge model.
-The GPT is prompted to judge whether the generated prediction semantically matches the ground truth answer.
-The prompt template is shown in Table 8.
-All the reported experimental results are from single runs.
+Table 9: Hyperparameters for each training process. Pre-Training and Fine-Tuning refer to the two stages of the standard MLLM training process.
+Appendix B Dataset Details
+Table 8 details the statistics of each dataset. Some datasets are filtered from their original ones:
+  • MSCOCO-2014 (Lin et al., 2014): Each image has multiple captions, we only use its first caption to form the training set.
+  • Clotho-AQA (Lipping et al., 2022): Each sample is annotated with a confidence level, we only use the samples whose confidence levels are "yes" to form the training set and test set.
+  • Cap3D (Luo et al., 2024): Since the original dataset is huge in scale, we filter out the samples whose caption is longer than 100 letters.
+Then, we randomly sample a 50K subset as the training set.
+  • Cap3D-QA (Panagopoulou et al., 2023): Since the original dataset is huge in scale, we randomly sample a 30K subset as the training set.
+For each dataset, we use a randomly sampled 1K subset of its holdout test set as the final test set, except for the MSVD (Chen and Dolan, 2011), since the size of its original test set is less than 1K.
+Appendix C Experimental Details
+C.1 Implementation Details
+We build our experimental codebase on top of LLaVA (Liu et al., 2024b, a) and NExT-GPT (Wu et al., 2024). We detail the modality-specific components of each modality as follows:
+  • Image: We use CLIP-ViT-L-336px (Openai, 2021) as the pre-trained image encoder, a randomly initialized MLP as the connector.
+  • Video: We use CLIP-ViT-L-336px (Openai, 2021) as the pre-trained video encoder, a randomly initialized MLP as the connector.
+We uniformly sample 4 frames from a video as input frames.
+Then each frame is encoded by the video encoder separately.
+The output feature frames are downsampled by 2x using bilinear pooling before sending into the connector to improve efficiency.
+  • Audio: We use $\mathrm{BEATs_{iter3+}(AS2M)}$ (Chen et al., 2023b) as the pre-trained audio encoder, a Q-Former (Li et al., 2023) initialized from the pre-trained bert-base-uncased (Kenton and Toutanova, 2019) as the connector.
+The number of query tokens is set to 32.
+  • Point Cloud: We use Point-BERT-v1.2 (Xu et al., 2024) as the pre-trained point cloud encoder, a randomly initialized MLP as the connector.
+We set the hyperparameters mainly following previous works (Liu et al., 2024b, a), as listed in Table 9.
+For training that involves updating the LLM backbone, we utilize parameter-efficient fine-tuning with LoRA (Hu et al., 2022) applied across all linear modules within the LLM, setting the LoRA rank to 128 and the alpha parameter to 128.
+All the experiments are conducted on a single NVIDIA RTX A6000 48G with FP16.
 System Prompt:
@@
 Table 10: Prompts to query the GPT for open-ended QA evaluation. The placeholders in red boxes are filled according to each evaluated sample.
-Appendix D Implementation of Baselines
+C.2 Evaluation Details
+For the calculation of CIDEr scores (Vedantam et al., 2015), we utilized an open-sourced library CaptionMetrics (wangleihitcs, 2019).
+For the calculation of prediction accuracy, we leverage a GPT-based open-ended QA evaluation with GPT-4o mini as the judge model.
+The GPT is prompted to judge whether the generated prediction semantically matches the ground truth answer.
+The prompt template is shown in Table 10.
+All the reported experimental results are from single runs.
+C.3 Implementation of Baselines
 The implementation details of each CL baseline are listed as follows:
-  • Replay leverage a replay dataset $R_{i}\leftarrow$ sample $r\%$ data from $\{D_{1},D_{2},\dots,D_{i-1}\}$ when training on a new modality $M_{i}$.
-It can be seen as training on a joint dataset of $R_{i}$ and $D_{i}$.
-  • EWC firstly estimates the Fisher information matrix $F_{i-1}$ of the last training stage $i-1$ as:
+  • Replay leverage a small replay dataset $R_{i}\leftarrow$ randomly sample rr% data from $\{D_{1},D_{2},\dots,D_{i-1}\}$.
+When training on a new modality $M_{i}$, it trains on the joint dataset of $R_{i}$ and $D_{i}$.
+  • EWC firstly estimates the Fisher information matrix $F_{i-1}$ of the last training stage i1i-1 as:
     $F_{i-1}=$
@@
   • PathWeave leverages Adapter-in-Adapter (AnA) modules.
-In our implementation, the AnA modules are injected in the LLM rather than the connector for a fair comparison.
+In our implementation, the AnA modules are injected into the LLM rather than the connector for a fair comparison.
 The rank of AnA is consistent with the LoRA rank of other baselines, which is 128.
-Appendix E Complete Raw Data
+In their original settings, PathWeave removes the newly added modules when testing the former modalities.
+However, this results in the inability to perform cross-modality tasks, which are common in real applications.
+Therefore, we do not remove them for a fair comparison.
+For each baseline, all the common parameters about training MLLM itself are the same and set as default in Table 9.
+Appendix D Complete Raw Data
 Table 11 and Table 12 show the raw data of Figure 3 and Figure 4.
@@
 Stage 4 56.54 0.227 36.72 0.564 26.64 0.651 96.40 0.551
+EProj Stage 1 92.16 0.298 - - - - - -
+Stage 2 92.16 0.298 123.04 0.470 - - - -
+Stage 3 92.16 0.298 123.04 0.470 54.85 0.637 - -
+Stage 4 92.16 0.298 123.04 0.470 54.85 0.637 58.59 0.349
 PathWeave Stage 1 100.76 0.358 - - - - - -
@@
 Stage 4 45.25 0.327 23.39 0.511 47.53 0.524 98.91 0.320
+EProj Stage 1 58.59 0.349 - - - - - -
+Stage 2 58.59 0.349 54.85 0.637 - - - -
+Stage 3 58.59 0.349 54.85 0.637 123.04 0.470 - -
+Stage 4 58.59 0.349 54.85 0.637 123.04 0.470 92.16 0.298
 PathWeave Stage 1 99.93 0.568 - - - - - -